### v1

# Perceptual Decoupling for Scalable Multi-modal Reasoning via Reward-Optimized Captioning

Abstract
Recent advances in slow-thinking language models (e.g., OpenAI-o1 and DeepSeek-R1) have demonstrated remarkable abilities in complex reasoning tasks by emulating human-like reflective cognition. However, extending such capabilities to the multi-modal large language models (MLLMs) remains challenging due to the high cost of retraining vision-language alignments when upgrading the underlying reasoner LLMs. A straightforward solution is to decouple perception from reasoning, i.e., converting visual inputs into language representations (e.g., captions) that are then passed to a powerful text-only reasoner. However, this decoupling introduces a critical challenge: the visual extractor must generate descriptions that are both faithful to the image and informative enough to support accurate downstream reasoning. To address this, we propose Reasoning-Aligned Perceptual Decoupling via Caption Reward Optimization (RACRO) —a reasoning-guided reinforcement learning strategy that aligns the extractor’s captioning behavior with the reasoning objective. By closing the perception-reasoning loop via reward-based optimization, RACRO significantly enhances visual grounding and extracts reasoning-optimized representations. Experiments on multi-modal math and science benchmarks show that the proposed RACRO method achieves state-of-the-art average performance while enabling superior scalability and plug-and-play adaptation to more advanced reasoning LLMs without the necessity for costly multi-modal re-alignment.
1 Introduction
Recent developments in slow-thinking language models, such as OpenAI-o1 [26], DeepSeek-R1 [22], Gemini-Thinking [11], and Qwen3 [45], have led to significant advancements in the performance of large language models (LLMs) for complex math and science tasks. These models replicate human-like reflective thinking, intentionally modeling reasoning as a step-by-step process rather than relying on superficial shortcuts. As a result, they surpass fast-thinking models like GPT-4o [25] and Claude 3.5 [2], showing improvements of over 30% on mathematical benchmarks such as AIME24 [1] and AMC23 [15], along with approximately 10% gains on science benchmarks like GPQA [41].
Motivated by these successes, there is growing interest in equipping multi-modal large language models (MLLMs) with similar slow-thinking capabilities to solve vision-based reasoning tasks that require both spatial perception and logical inference. A common strategy—adopted by methods like VL-Rethinker [47] and MM-EUREKA [38]—is to fine-tune a pre-trained MLLM (e.g., Qwen2.5-VL [5]) on multi-modal reasoning datasets using reinforcement learning with verifiable reward signals. However, this approach faces two critical limitations: (1) it is inherently restricted by the capabilities of underlying LLMs used for initialization; (2) adapting to more advanced LLMs necessitates redoing the vision-language alignment process, which involves burdensome multi-modal training that requires computing trillions of tokens. This raises a challenging research question: Can we efficiently integrate slow-thinking strategies into multi-modal reasoning without retraining the entire multi-modal LLM?
A seemingly intuitive solution is to utilize a perceptual decoupling approach: separate the vision and reasoning components by converting multi-modal inputs into natural language captions, which are then passed to an external, more powerful text-only reasoner. While this idea promises flexibility and modularity, it introduces a fundamental challenge—the success of the reasoning process now critically depends on the quality of intermediate captions. If the visual extractor generates captions that are incomplete, overly generic, or hallucinatory, even the strongest reasoning LLMs will fail to give correct reasoning processes. As shown in Figure 3, existing MLLMs might struggle to produce reasoning-relevant descriptions, especially for complex mathematical and scientific visual inputs [51].
Figure 1: Scalable multi-modal reasoning of RACRO. RACRO reveals better scalability compared to the traditional visual alignment, while remaining computationally efficient.
To overcome this limitation, we propose Reasoning-Aligned Perceptual Decoupling via Caption Reward Optimization (RACRO) —a unified framework that not only decouples visual perception from reasoning, but also ensures their alignment through a reasoning-aware reinforcement learning procedure. In the proposed RACRO method, an MLLM-based extractor is responsible for encoding task-relevant visual information to natural language, while an external text-only LLM reasoner handles the slow-thinking reasoning process. Crucially, we introduce Caption Reward Optimization (CRO), a reinforcement learning algorithm that fine-tunes the extractor using reasoning outcomes as feedback. Given an input image and a question, the extractor generates multiple caption candidates; each is passed to the reasoner, and rewards are assigned based on whether the reasoner generates correct answers or not. This feedback loop aligns the extractor with the reasoning objective, guiding it to generate faithful and problem-relevant descriptions.
By integrating perceptual decoupling with CRO into a unified framework, we enable effective and scalable slow-thinking in multi-modal reasoning systems. Experiments on challenging benchmarks—such as MathVerse [56], MathVision [48] and LogicVista [52]—show that the proposed RACRO method achieves state-of-the-art results. Furthermore, the proposed RACRO method supports plug-and-play adaptation to novel reasoning LLMs without extra multi-modal retraining, making it a practical and scalable solution for the rapid evolution of foundation models. It is illustrated in Figure 1, comparing the average performance of various MLLMs and their RACRO-enhanced variants across model sizes. For each RACRO-enhanced model group (e.g., Qwen2.5-VL-3B ($\bf{RACRO}$)), the original variant (leftmost, solid) is only trained with a base LLM reasoner with a small amount of data (e.g., 39K) using CRO, which can then achieve scalable performance with stronger LLMs (see detailed configurations on the choices of LLMs in Appendix A). Connecting these points yields a novel Pareto front that captures a superior trade-off between model size and performance than other base MLLMs.
Our contributions can be summarized as the following three perspectives:
  • We propose RACRO, a unified framework that decouples visual perception from reasoning and aligns them through reinforcement learning with verifiable rewards (RLVR).
  • We introduce Caption Reward Optimization, a novel reinforcement learning strategy that uses reasoning success as a reward signal to optimize the extractor’s captioning behavior.
  • We obtain state-of-the-art performance on multi-modal reasoning benchmarks, while maintaining scalability and eliminating the necessity for costly vision-language re-alignment.
(a) Existing MLLM alignment methods
(b) RACRO (ours)
Figure 2: Comparisons between existing alignment methods for reasoning MLLMs and RACRO. For novel LLMs, existing methods (a) repeatedly conduct the intensive alignment process, while (b) RACRO decouples visual perception from reasoning (Sec. 3.1) by learning to extract reasoning-aligned visual contexts with the proposed CRO paradigm (Sec. 3.2).
2 Related Work
Multi-modal Large Language Models (MLLM).
With the development of LLMs, integrating visual perception capability into LLMs has received increasing attention [19, 27, 30]. To address the gap between different modalities, existing paradigms [5, 20, 21, 59] start with a text-only LLM and conduct pre-training on large-scale image captioning data to obtain a vision-language connector. Then, supervised fine-tuning (SFT) is applied with visual instruction tuning datasets. It is important to note that this procedure is computationally intensive. For example, Qwen2.5-VL [5] consumes around 4.1T tokens, and Kimi-VL [44], a more recent MLLM with MoE architecture, requires 2.3T tokens for the alignment. Under this paradigm, it remains formidable to upgrade MLLMs to newer LLMs. Instead, this paper proposes a computationally efficient solution for multi-modal reasoning.
Empowering MLLMs with Slow-Thinking Capabilities.
To empower MLLM with similar slow-thinking abilities, a widely adopted approach is to start with an existing MLLM (e.g., Qwen2.5-VL [5]) and perform reinforcement learning or knowledge distillation. For example, VL-Rethinker [47], MM-EUREKA [38] and NoisyRollout [31] apply GRPO [42] (or its variants [13, 31, 32, 38]) to MLLMs to learn deliberate reasoning patterns. Distillation-based methods, such as the OpenVLThinker [13], R1-OneVision [54], and Vision-R1 [24], perform supervised fine-tuning (SFT) on the vision-language reasoning data generated via a caption-and-reasoning pipeline [19], where visual inputs are converted to textual descriptions first, and strong text-only LLMs are then utilized to generate reasoning traces. However, their performances are limited by the base LLM (e.g., Qwen2.5 [53]), which often lags behind state-of-the-art reasoning models (e.g., DeepSeek-R1-Distilled [22]) in tasks like math. While replacing it with a stronger LLM is an intuitive solution, re-aligning vision and language through full retraining on trillions of tokens is prohibitively costly. In this work, we introduce the RACRO, a reasoning-aligned perceptual decoupling framework that bridges perception and reasoning via reward optimization, allowing seamless integration of stronger LLMs without retraining the vision-language interface, and meanwhile, demonstrating superior scalability w.r.t model sizes (c.f., Figure 1).
3 Methodology
In this section, we propose Reasoning-Aligned Caption Reward Optimization (RACRO), a unified framework that enables multi-modal slow-thinking reasoning by aligning the visual perception module with the downstream reasoning objectives. RACRO builds on the intuition of perceptual decoupling, where visual inputs are first translated into natural language representations that can be processed by powerful text-only reasoners. However, rather than treating the decoupling as static and assuming fixed captions suffice, RACRO closes the perception-reasoning loop through reinforcement learning, optimizing the captioning behavior of the visual extractor for downstream reasoning success.
3.1 Reasoning-Aligned Perceptual Decoupling
The perceptual decoupling paradigm is based on the assumption that the process of solving visual reasoning tasks can be separated into two distinct stages: visual perception and text-only reasoning. Specifically, given an image and a related question, if all relevant visual details can be losslessly converted into natural language, the problem can be solved using a text-only LLM. This decoupling allows for seamless upgrades to more powerful LLMs without the need to retrain or realign vision-language models. However, extracting all necessary visual details remains a severe challenge, as visual information can be dense, complex, and ambiguous [19, 51]. To address that, we propose the Reasoning-Aligned Perceptual Decoupling framework, focusing on providing reasoning-aware visual details to the LLM instead of attempting to capture the entire visual input. Figure 2 gives an overview of the framework. Specifically, It consists of two steps:
Figure 3: Inaccurate visual captions result in reasoning failure. MLLMs might produce hallucinated (left), vague, or incomplete captions (middle), particularly for complex visual reasoning, while captions containing necessary query-relevant contexts are essential for successful reasoning (right).
Extraction.
Guided by the goal of reasoning-aligned perceptual decoupling, we design an extraction process that generates visual details tailored to the reasoning needs for a given question. Instead of producing generic descriptions, we leverage an MLLM, such as Qwen2.5-VL [5] or InternVL3 [59], denoted as $M_{V}$, to produce two complementary outputs from an image $I$ and a query $\mathbf{q}$: a query-conditioned caption $\mathbf{c}=M_{V}(I,P_{\text{cap}}(\mathbf{q}))$, and a tentative solution $\hat{\mathbf{y}}=M_{V}(I,P_{\text{cot}}(\mathbf{q}))$, with each generated using distinct prompts (see Appendix B). The caption $\mathbf{c}$ captures visual content which is selectively relevant to the question, functioning as a reasoning-aligned abstraction. Although $\hat{\mathbf{y}}$ may not be fully accurate due to the limited reasoning capabilities of MLLMs, it often contains useful inductive signals that can inform and guide further downstream reasoning inference.
Reasoning.
The reasoning component is separately handled by a text-only large language model (LLM) (e.g., R1-Distilled-7B/32B [22], Qwen3-8B [45], or QwQ-32B [46]), denoted as $M_{R}$. It takes as input the question $\mathbf{q}$, the reasoning-aligned caption $\mathbf{c}$, and the tentative solution $\hat{\mathbf{y}}$, and produces the final answer: $\mathbf{y}=M_{R}(P_{\text{reason}}(\mathbf{q},\mathbf{c},\hat{\mathbf{y}}))$, where $P_{\text{reason}}$ is a prompt formatting function that integrates the inputs into a coherent textual reasoning context (see Appendix B).
A core advantage of this decoupled framework is its flexibility: by separating reasoning from visual perception, we can boost performance on multi-modal tasks simply by upgrading to more capable LLMs for $M_{R}$, without modifying or retraining the MLLMs. While concurrent works have adopted similar decoupling paradigms—either as data engines for generating reasoning traces [24, 54] or as the inference framework [37]—they fail to investigate how to generate 1) query-conditioned and 2) reasoning-aligned visual representations, which is the focus of this paper in the following.
3.2 Caption Reward Optimization
While this design enables flexible upgrade of the reasoner without vision-language re-alignment, it exposes the system to a critical bottleneck: the quality of the caption directly affects the reasoning outcome. In practice, existing MLLMs may generate hallucinated [51], vague, or incomplete captions, particularly for complex visual reasoning tasks such as the geometric math questions (see Figure 3). To address this bottleneck, RACRO introduces Caption Reward Optimization—a reasoning-guided reinforcement learning algorithm that fine-tunes the captioning extractor using the feedback from a text-only reasoner, aiming at reinforcing captions that maximize likelihood of successful reasoning.
Caption Reward Optimization
is inspired by GRPO (Group Relative Policy Optimization) [42], which is a policy optimization algorithm originally developed to enhance the reasoning capability of text-only LLMs. In our setting, the policy $\pi_{\theta}$ to optimize becomes the MLLM $M_{V}$ performing visual captioning. For a given input pair $(I,\mathbf{q})$ of image and text question from the training set $p_{\mathcal{D}}$, the old policy generates $G$ caption rollouts, i.e., $\mathbf{c}\sim$ $\pi_{\theta_{\mathrm{old}}}(I,P_{\text{cap}}(\mathbf{q}))$. Denoting $R_{i}$ as the reward for the $i$-th rollout, the normalized advantage is $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ denotes the standard deviation of rewards within the group and the baseline reward is $\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$. The objective incorporates a surrogate loss clipped within $[1-\epsilon,1+\epsilon](\epsilon>0)$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown here) to stabilize optimization:
$L(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\mathbf{c}\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{cap}}(\mathbf{q}))}$
$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl{(}\frac{\pi_{\theta}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl{(}\frac{\pi_{\theta}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))},\,1-\epsilon,\,1+\epsilon\Bigr{)}\hat{A}_{i}\Biggr{)}\Biggr{]}\textrm{.}$ (1)
However, different from the math reasoning problem traditionally studied by GRPO, there are no extractable answers in the generated captions. Therefore, the widely adopted rule-based reward is no longer applicable to our problem. To tackle this, we propose to combine the rule-based reward function $r(\cdot,\cdot)$ and a reasoning LLM $M_{R}$ for reward computation. Specifically, the reward $R_{i}$ of a caption rollout $\textbf{c}_{i}$ is determined by whether a reasoner LLM $M_{R}$ could arrive at the ground-truth answer using this caption $\textbf{c}_{i}$ as contexts, which can be expressed as:
$R_{i}=r(y_{\text{gt}},y_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(y_{i})),\quad\text{where}\ y_{i}=M_{R}(P_{\text{reason}}^{\prime}(\textbf{q},\textbf{c}_{i})),$ (2)
where $y_{\text{gt}}$ denotes the ground-truth answer of a reasoning question and $\mathbbm{1}(\cdot)$ is an indicator function that outputs 1 if the final parsed prediction matches the ground-truth and 0 otherwise. Note that the reasoning process during reward computation is different from that in Sec. 3.1 as the tentative response is not provided here (check the ablations on this configuration in Table 3). In summary, the proposed Caption Reward Optimization offers the following benefits:
  • More accurate and task-relevant captions: As defined in Eq. (2), the reward encourages captions that facilitates successful downstream reasoning. Consequently, the MLLM learns to generate captions with less visual hallucination and more task-relevant visual details.
  • Alignment scalability to any reasoners: Once the MLLM undergoes the Caption Reward Optimization process, even LLMs that are not involved during training can be incorporated into the reasoning framework because the captions are natural languages that generalize between different LLMs. This eliminates repeated alignments as done in the traditional MLLM instruction tuning approaches. As demonstrated in Figure. 1, our RACRO reveals superior scalability during inference time, even though not been specifically trained for.
Figure 4: Caption Reward Optimization reinforces captions that induce correct reasoning results via reinforcement learning with verifiable rewards (RLVR).
4 Experiments
4.1 Main Results
Evaluation.
Following VisualPRM [49], we evaluate our model on a diverse set of multi-modal reasoning benchmarks, including the MathVista (testmini) [36], MathVision (test) [48], MathVerse (vision-only) [56], MMMU (val) [55], WeMath [39], DynaMath [60], and the LogicVista [52]. We report the worst-case accuracy for DynaMath and the strict accuracy for WeMath, while for the other benchmarks, we report the overall accuracy. All evaluations are conducted using VLMEvalKit [14].
Table 1: Comparison on multi-modal reasoning benchmarks.
Model Math Vista Math Vision Math Verse MMMU We Math Dyna Math Logic Vista AVG
Proprietary Models
Claude-3.7-Sonnet 66.8 41.9 46.7 75.0 49.3 39.7 58.2 53.9
Gemini-2.0-Flash 70.4 43.6 47.7 72.6 47.4 42.1 52.3 53.7
GPT-4o-20241120 60.0 31.2 40.6 70.7 45.8 34.5 52.8 47.9
o1 73.9 42.2 - 78.2 - - - -
Open-Source Models
InternVL3-2B-MPO 57.0 21.7 25.3 48.6 22.4 14.6 36.9 32.4
OpenVLThinker-7B 70.2 25.3 - - - - - -
R1-Onevision-7B 64.1 29.9 40.0 - - - - -
Adora-7B 73.5 23.0 - - - - - -
MM-Eureka-7B 73.0 26.9 - - - - - -
InternVL3-8B-MPO 71.6 29.3 39.8 62.7 37.1 25.5 44.1 44.3
InternVL3-14B-MPO 75.1 37.2 44.4 67.1 43.0 31.3 51.2 49.9
MM-Eureka-32B 74.8 34.4 - - - - - -
InternVL3-38B-MPO 75.1 34.2 48.2 70.1 48.6 35.3 58.4 52.8
LLaVA-OV-72B 67.5 30.1 39.1 56.8 - - - -
EMOVA-72B 69.9 - 50.0 59.7 - - - -
VL-Rethinker-72B 80.3 43.9 - 68.8 - - - -
Qwen2.5-VL-72B 74.2 39.3 47.3 68.2 49.1 35.9 55.7 52.8
QVQ-72B-Preview 70.3 34.9 48.2 70.3 39.0 30.7 58.2 50.2
InternVL3-78B-MPO 79.0 43.1 51.0 72.2 46.0 35.1 55.9 54.6
Qwen2.5-VL-3B 64.5 21.9 28.8 50.1 24.2 13.4 39.6 34.6
w/ VisualPRM-8B-v1.1 69.5 35.1 36.3 56.0 39.2 10.6 43.4 41.4
w/ RACRO (R1-7B) 64.5 (+0.0) 36.3 (+14.4) 39.3 (+10.5) 51.7 (+1.6) 31.5 (+7.3) 21.4 (+8.0) 46.5 (+6.9) 41.6 (+7.0)
w/ RACRO (Qwen3-8B) 68.6 (+4.1) 39.6 (+17.7) 43.7 (+14.9) 58.6 (+8.5) 32.3 (+8.1) 29.3 (+15.9) 51.9 (+12.3) 46.3 (+11.7)
w/ RACRO (R1-32B) 70.2 (+5.7) 41.5 (+19.6) 43.8 (+15.0) 58.4 (+8.3) 32.5 (+8.3) 30.3 (+16.9) 55.7 (+16.1) 47.5 (+12.9)
Qwen2.5-VL-7B 70.3 25.8 41.0 57.3 34.4 19.4 46.1 42.0
w/ VisualPRM-8B-v1.1 74.6 27.1 45.0 56.1 46.3 17.8 48.5 45.1
w/ RACRO (R1-7B) 71.9 (+1.6) 38.9 (+13.1) 45.3 (+4.3) 58.7 (+1.4) 37.1 (+2.7) 27.0 (+7.6) 49.7 (+3.6) 46.9 (+4.9)
w/ RACRO (Qwen3-8B) 70.0 (-0.3) 41.5 (+15.7) 44.8 (+3.8) 60.6 (+3.3) 38.6 (+4.2) 26.5 (+7.1) 50.1 (+4.0) 47.4 (+5.4)
w/ RACRO (R1-32B) 74.4 (+4.1) 42.5 (+16.7) 46.6 (+5.6) 61.1 (+3.8) 31.0 (-3.4) 30.3 (+10.9) 55.0 (+8.9) 48.7 (+6.7)
Qwen2.5-VL-32B 76.8 37.8 50.1 69.0 43.1 33.3 55.0 52.2
w/ VisualPRM-8B-v1.1 77.7 39.9 52.2 69.6 50.4 33.1 53.9 53.8
w/ RACRO (R1-7B) 75.4 (-1.4) 44.5 (+6.7) 52.9 (+2.8) 68.8 (-0.2) 48.3 (+5.2) 35.1 (+1.8) 54.6 (-0.4) 54.2 (+2.0)
w/ RACRO (Qwen3-8B) 75.5 (-1.3) 47.1 (+9.3) 54.9 (+4.8) 67.9 (-1.1) 50.0 (+6.9) 35.7 (+2.4) 58.2 (+3.2) 55.6 (+3.4)
w/ RACRO (R1-32B) 75.9 (-0.9) 46.7 (+8.9) 55.6 (+5.5) 69.0 (+0.0) 48.9 (+5.8) 37.1 (+3.8) 58.6 (+3.6) 56.0 (+3.8)
Baselines.
We compare our method with the following baselines:
  • Proprietary models: GPT-4o [25], Claude-3.7-Sonnet [3], and Gemini-2.0-Flash [12] are general-purpose multi-modal models without dedicated reasoning optimization. In contrast, OpenAI o1 [26] is explicitly designed to enhance reasoning capabilities. These models represent the current proprietary state-of-the-art.
  • Open-source general-purpose models: LLaVA-OneVision-72B (LLaVA-OV-72B) [27], EMOVA-72B [6], and the Qwen2.5-VL series [5] are primarily pretrained or instruction-tuned on large-scale image-text datasets, without explicit reasoning alignment.
  • Open-source reasoning models: The InternVL3-MPO variants [59], OpenVLThinker-7B [13], R1-Onevision-7B [54], Adora-7B, MM-Eureka (7B/32B) [38], VL-Rethinker (7B/72B) [47], and QVQ-72B-Preview [40]. These models are explicitly fine-tuned for reasoning tasks using methods such as SFT, DPO, and rule-based reinforcement learning.
  • Multi-modal test-time scaling methods: We also compare our proposed method with multi-modal test-time scaling methods such as Best-of-N search (BoN) [49], which has been shown to effectively improve the reasoning accuracy of MLLMs. In particular, we consider the latest version, VisualPRM-8B-v1.1 [49]. Note that in Table 1, BoN is adopted independently with RACRO.
Implementation.
To validate the effectiveness of the proposed method, we perform RACRO on the following MLLMs: Qwen2.5-VL series (3B, 7B, and 32B). During training, we use R1-Distilled-7B (referred to interchangeably as R1-7B) as the reasoner to compute reward signals for all MLLMs. During evaluation, in addition to R1-Distilled-7B, we also employ its 32B counterpart and Qwen3-8B to assess the scalability of our method. For training data, we adopt ViRL39K [47], a curated dataset of 38,870 verifiable multi-modal question-answer pairs tailored for reinforcement learning. We implement the caption reward optimization algorithm using the verl [43] framework with a global batch size of 256, a rollout temperature of 1.0, and a learning rate of $1\mathrm{e}^{-6}$. For each query, four captions are sampled. The clipping constant $\epsilon$ and KL-penalty coefficient $\beta$ are set to 0.2 and $1\mathrm{e}^{-3}$.
Results.
Table 1 presents the performance of the proposed method. As shown, the average scores across seven datasets are significantly improved compared to the original MLLMs. For example, applying RACRO to Qwen2.5-VL-7B with a same-sized reasoner (R1-Distilled-7B) yields a +4.9% gain. Further improvements are observed when using a larger reasoner, such as a +6.7% increase with R1-Distilled-32B. Notably, when applying RACRO to Qwen2.5-VL-32B and evaluating with R1-Distilled-32B, we achieve state-of-the-art results on MathVision, MathVerse, and LogicVista, outperforming even larger models such as InternVL3-78B-MPO. Importantly, the Qwen3-8B and R1-32B reasoners are only used during inference time via a plug-and-play manner, since all trainings are conducted solely with the R1-7B reasoner. This demonstrates the scalability and generalization ability of our RACRO across both model sizes and (heterogeneous) reasoning LLM variants.
Compared to Best-of-8 (Bo8) search, RACRO also achieves superior results. Notably, for stronger MLLMs such as Qwen2.5-VL-32B, VisualPRM-8B-v1.1 provides only limited improvement (e.g., +1.6%). RACRO, on the other hand, demonstrates greater scalability to stronger reasoners: when paired with a same-sized reasoner as VisualPRM-8B-v1.1, it improves the performance of Qwen2.5-VL-32B by +3.4%. This improvement is achieved without any re-training, highlighting RACRO’s plug-and-play compatibility with more capable reasoners.
Table 2: Ablation study of different components of our RACRO.
Sol. Cap. LLM CRO Math Vista Math Vision Math Verse MMMU We Math Dyna Math Logic Vista AVG
31.5 36.0 10.0 39.2 11.4 3.8 22.8 22.1
Qwen2.5-VL-3B
64.5 21.9 28.8 50.1 24.2 13.4 39.6 34.6
60.1 36.3 30.2 47.6 24.4 14.6 44.5 36.8
59.5 36.2 12.7 47.8 25.0 13.4 35.1 32.8
63.8 34.5 30.3 51.7 27.4 17.4 43.9 38.4
64.5 36.3 39.3 51.7 31.5 21.4 46.5 41.6
Qwen2.5-VL-7B
70.3 25.8 41.0 57.3 34.4 19.4 46.1 42.0
68.3 38.0 41.0 53.9 31.2 22.4 47.4 43.2
54.4 35.8 27.9 50.6 23.7 7.6 34.9 33.6
70.9 37.6 42.3 56.4 37.3 24.6 48.3 45.3
71.9 38.9 45.3 58.7 37.1 27.0 49.7 46.9
4.2 Ablation Study
In this section, we first investigate the effectiveness of the proposed components, i.e., Reasoning-Aligned Perceptual Decoupling and Caption Reward Optimization. Then, we ablate the choices on reward computation during training. We use the same training configurations as in Sec. 4.1. Qwen2.5-VL-3B/7B are adopted for ablations due to resource constraints. Unless otherwise specified, we use R1-Distilled-7B as the default LLM reasoner throughout the training and inference stages.
Reasoning-Aligned Perceptual Decoupling & Caption Reward Optimization.
For simplicity, we use the following shorthand markers to indicate the components:
  • LLM indicates the usage of the text-only LLM reasoner $M_{R}$.
  • Cap. denotes the use of the query-conditioned captions generated by the MLLM $M_{V}$.
  • Sol. denotes the (tentative) solution provided by the MLLM $M_{V}$. Note that if the reasoner $M_{R}$ (i.e., LLM above) is not used, it is considered as the final output solution.
  • CRO denotes the use of Caption Reward Optimization that optimizes the quality of the generated captions.
We assess the RACRO framework by progressively enabling its components, as in Table 2. Starting from the baseline MLLM, we first introduce the tentative solution as the input to the LLM reasoner. This alone leads to notable performance improvements—e.g., from 34.6% to 36.8% for Qwen2.5-VL-3B and from 42.0% to 43.2% for the 7B model—demonstrating the reasoner’s ability to refine intermediate outputs from the MLLM. However, on tasks such as MathVista, MMMU, and WeMath, performance declines after introducing the LLM with tentative solutions only. To compensate for the lack of visual details, we incorporate query-conditioned captions alongside the tentative solution, which effectively restores the performance. Surprisingly, when captions are adopted alone without tentative solutions, benchmark performance drops significantly, suggesting that while captions can provide valuable visual context, they are often noisy—suffering from hallucinations and missing critical details. In contrast, the tentative solution acts as a reliable fallback for the reasoner when the visual signal is weak or ambiguous. To improve the quality of the query-conditioned captions, we apply the proposed Caption Reward Optimization strategy. As a result, we observe consistent performance gains across all tasks, confirming the quality issue of the captions and the effectiveness of the Caption Reward Optimization strategy.
Table 3: Ablation studies on (a) reasoner types and (b) inputs for reward computation.
Configuration Q-VL-3B Q-VL-7B
w/o CRO (baseline) 38.4 45.3
(a) Reasoner Types
Qwen-2.5-7B-Instruct 40.4 45.1
R1-Distilled-1.5B 40.9 45.2
R1-Distilled-7B 41.6 46.9
(b) Reasoner Inputs
Cap. & Sol. 40.8 45.8
Cap. 41.6 46.9
Choices on Reward Computation.
Conducting CRO requires computing the reward of a caption rollout, which involves running the reasoner LLM. Here we discuss (i) what type of reasoner and (ii) what kind of inputs to the reasoner lead to more accurate reward computation. Specifically, for reasoners, we consider R1-Distilled-1.5B/7B and Qwen-2.5-7B-Instruct. For the reasoner input, we consider “Cap.”, which only includes the generated caption and “Cap. & Sol.” that further includes the MLLM’s tentative solution (pre-generated before training). Note that the later is exactly the configuration adopted in the proposed Perceptual Decouple framework.
Table 3 shows the average performance with Qwen2.5-VL series as MLLMs (i.e., Q-VL-3B/7B) under different reasoner types and inputs. We observe that using smaller reasoners (i.e., R1-Distilled-1.5B) or models less specialized in reasoning (i.e., Qwen-2.5-Instruct) leads to worse overall performance. Intuitively, models performing poorly in reasoning are prune to generate incorrect solutions even though captions are accurate and comprehensive, providing noisy reward signals for CRO. However, larger reasoners significantly slow down the reward computing. Thus, we adopt R1-Distilled-7B as the reasoner as it strikes a balance between computational efficiency and reward accuracy by default.
Regarding the input to the reasoner, we found passing only captions yields the best performance. Although “Cap. & Sol.” perfectly aligns with the Perceptual Decoupling framework during inference time, including a tentative response, allows the reasoner to take a shortcut that ignore the captions during reasoning, thereby producing a noisy reward signal for the captions.
4.3 Analysis
Scalability of CRO.
In a practical scenario, after conducting CRO on an MLLM with a certain LLM (e.g., R1-Distilled-7B), it is desirable to switch to new LLMs (e.g., QwQ-32B) without additional training. We refer to this as the scalability of CRO and reveal that CRO indeed possesses this property.
Figure 6 shows the overall task performance under the Perceptual Decoupling framework using different reasoners across varying model sizes and generations. The MLLMs adopted are from the Qwen2.5-VL series, both with and without CRO. As shown, although all MLLMs were optimized with CRO using R1-Distilled-7B, switching to different LLMs—whether in size (R1-Distilled-1.5B/32B) or generation (QwQ-32B)—still results in a notable performance boost over their non-CRO counterparts. This confirms the scalability of CRO. Furthermore, we observe that this performance gain becomes more pronounced with larger or stronger LLMs. This is because, without CRO, these models are significantly bottlenecked by inaccurate or low-quality visual captions, preventing them from leveraging their full reasoning potential. In contrast, smaller or weaker LLMs are limited not only by caption quality but also by their own reasoning capacity, making the relative improvement from CRO less substantial. This highlights the effectiveness of CRO in improving the quality of image captions. Further, with stronger LLMs, the performance gap between 3B and 7B gradually diminishes. This reveals that reasoning is better decoupled from perception.
Figure 5: Ablation on the choice of reasoner.
Figure 6: Ablation on CRO and GRPO.
CRO and GRPO are Complementary.
The proposed CRO objective effectively enhances the quality of query-conditioned captions, thereby improving overall performance. Another direction for improving the reasoning capability of the LLM is to produce more accurate tentative responses, which can be achieved through reinforcement learning methods such as GRPO. To evaluate their contributions, we apply CRO and GRPO separately to the original MLLMs (e.g., Qwen2.5-VL) and compare their average task performance under the perceptual decoupling framework (details on GRPO are in Appendix C.1). To investigate whether the two objectives are complementary, we also assess a combined setting—denoted as “CRO+GRPO”—where tentative responses and captions are generated by models fine-tuned separately with GRPO and CRO. For comparison, results from models without additional training are labeled as “No Training”. Additionally, we further include VL-Rethinker-7B [47], which is initialized from Qwen2.5-VL-7B and thoroughly fine-tuned using GRPO. We further fine-tune this model with CRO and report the resulting performance as “GRPO+CRO”. All results are visualized in Figure 6.
As shown, both CRO and GRPO improve the performance under the perceptual decoupling framework. More importantly, for all MLLMs, combining both objectives brings notable improvement over the individuals. This suggests that the proposed CRO is complementary to GRPO. Note that we did not report results with GRPO training in Table 1 to highlight the impact of perceptual decoupling.
Figure 7: Pairwise comparison on caption quality.
Evaluating the Quality of Captions.
To see whether the quality of the generated captions are improved, we conduct pairwise comparisons [9, 34] among captions generated by Qwen2.5-VL-3B trained with and without CRO. We randomly sample 1000 captions for each evaluation dataset and use GPT-4o  [^1] https://genai.ust.hk/ as the judge to determine the preferred caption. The judge receives the image, the question alongside two captions, and decides the winner or a tie game. We instruct the judge to choose captions that include more comprehensive and accurate details required to answer the question and exclude any solving process in the captions (the detailed prompt is in Appendix C.2). Furthermore, we alternate the order of the two captions to avoid position bias following [57]. Results are reported in Figure 7. It can be observed that across all datasets, models trained with CRO offer a clear advantage over the original ones. This confirms the quality issues of captions discovered by this paper and highlights the effectiveness of CRO. To ensure robustness of our automated evaluation, we further conduct human assessment through the same pairwise comparison framework with details in Appendix C.2. The results reinforces the quality improvement of the captions.
5 Conclusion
In this paper, we propose the RACRO framework to construct multi-modal reasoning models efficiently. By decoupling multi-modal reasoning with the visual perception and text-only reasoning, RACRO embraces the extraordinary reasoning capabilities of frontier LLMs, while getting rid of the burdensome visual re-alignment. Enhanced with Caption Reward Optimization, precise captions are reinforced to provide abundant visual contexts for better downstream reasoning, effectively scaling towards more advanced LLM reasoners during inference time. We obtain state-of-the-art performance on multiple multi-modal reasoning benchmarks, while remaining computationally efficient.
Appendix
Appendix A Model Configuration for Figure 1
For each RACRO-enhanced model group (e.g., Qwen2.5-VL-3B (RACRO)) in Figure 1, we train the original model (e.g., Qwen2.5-VL-3B, they are shown as hollow squares with different colors in the figure) using both the proposed CRO and the existing GRPO objectives, where the former encourages the MLLM to generate query-conditioned captions with higher quality while the later optimizes it to give better reasoning traces. The details for CRO and GRPO can be found in Sec. 4.1 and Appendix C.1. We then pair the trained MLLM with different LLM reasoners under our RACRO framework. Table 4 shows the configuration for each RACRO-enhanced model group. Specifically, for each model in the group, we report the total model size (B), paired LLMs, data used to conduct GRPO and CRO and the average performance across 7 tasks. Note that the results for Qwen2.5-VL-32B group does not involve GRPO training because we obtain worse results than the original model after GRPO. Results for other MLLMs without RACRO are directly taken from Table 1.
Table 4: Model Configurations for each RACRO-enhanced model group in Figure 1. “-” denotes the corresponding item (e.g., LLM reasoner, GRPO/CRO training) is not applied.
Size (B) LLM GRPO CRO Avg. Performance
Qwen2.5-VL-3B
3 - - - 34.6
7 Qwen3-4B ViRL39K ViRL39K 45.4
11 Qwen3-8B ViRL39K ViRL39K 46.3
35 R1-32B ViRL39K ViRL39K 47.8
Qwen2.5-VL-7B
7 - - - 44.5
11 Qwen3-4B ViRL39K ViRL39K 48.9
15 Qwen3-8B ViRL39K ViRL39K 50.2
39 R1-32B ViRL39K ViRL39K 50.8
Qwen2.5-VL-32B
32 - - - 52.2
39 R1-7B - ViRL39K 54.2
40 Qwen3-8B - ViRL39K 55.6
64 R1-32B - ViRL39K 56.0
Appendix B Prompt Templates Used in RACRO
In Table 5, we provide an index of the prompt templates used in RACRO.
Table 5: Index of Prompt templates for RACRO.
Component Purpose Notation Prompt Template
MLLM Tentative response $P_{\text{cot}}$ Figure 10
MLLM Query-conditioned captions $P_{\text{cap}}$ Figure 11
LLM Reasoner Inference $P_{\text{reason}}$ Figure 8
LLM Reasoner Reward computation $P^{\prime}_{\text{reason}}$ Figure 9
Figure 8: Prompt templates used by the reasoner LLM for inference.
Figure 9: Prompt templates used by the reasoner LLM for training.
Figure 10: Prompt templates used by the MLLM to obtain the tentative response. The placeholder is for the question.
Figure 11: Prompt templates used by the MLLM to obtain the query-conditioned captions.
Appendix C Experimental Details
C.1 Details for GRPO Training
GRPO
(Group Relative Policy Optimization) [42] is a policy optimization algorithm originally developed to enhance the reasoning capability of text-only LLMs. In our setting, the policy $\pi_{\theta}$ to optimize becomes the MLLM. For a given input pair $(I,\mathbf{q})$ of image and text question from the training set $p_{\mathcal{D}}$, the old policy generates $G$ rollouts, i.e., $\mathbf{o}\sim$ $\pi_{\theta_{\mathrm{old}}}(I,P_{\text{cot}}(\mathbf{q}))$. Denoting $R_{i}$ as the reward for the $i$-th rollout, the normalized advantage is $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ denotes the standard deviation of rewards within the group and the baseline reward is $\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$. The objective incorporates a surrogate loss clipped within $[1-\epsilon,1+\epsilon](\epsilon>0)$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown here) to stabilize optimization:
$L(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\mathbf{o}\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{cot}}(\mathbf{q}))}$
$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))},\,1-\epsilon,\,1+\epsilon\Bigr{)}\hat{A}_{i}\Biggr{)}\Biggr{]}\textrm{.}$ (3)
The reward $R_{i}$ for the $i$-th rollout is expressed as: $R_{i}=r(y_{\text{gt}},\textbf{o}_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(\textbf{o}_{i})),$ where $y_{\text{gt}}$ denotes the ground-truth answer of a reasoning question and $\mathbbm{1}(\cdot)$ is an indicator function that outputs 1 if the final parsed prediction matches the ground-truth and 0 otherwise.
Implementation Details.
Similar to CRO, for training data, we adopt ViRL39K [47] and implement GRPO using the verl [43] framework with a global batch size of 256, a rollout temperature of 1.0, and a learning rate of $1\mathrm{e}^{-6}$. For each query, eight rollouts are sampled. The clipping constant $\epsilon$ and KL-penalty coefficient $\beta$ are set to 0.2 and $1\mathrm{e}^{-3}$, respectively.
C.2 Details for Pairwise Comparisons
Prompt for GPT evaluations.
We provide the prompt for GPT evaluations on the quality of the caption in Figure 12. Consistent with the details in Sec. 4.3, this instructs the GPT to (1) choose captions that include more comprehensive and accurate details required to answer the question and (2) exclude any solving process in the captions.
Figure 12: Prompt templates used for GPT evaluations on caption qualities.
Human evaluations.
We present the details of the human evaluation conducted for the pairwise comparison experiment. For this, 100 questions are randomly sampled from the testmini set of MathVista, and captions are generated using Qwen2.5-VL-3B, trained with and without CRO. A total of 4 trained human annotators are recruited, with each annotator responsible for evaluating 25 questions to determine a winner or a tie.
The annotators’ primary task is to evaluate a question and its two anonymous model-generated answers, then select the one they deem of higher quality, following the scoring principle adopted from [4]. If neither answer stood out, they are instructed to choose "tie."
In Table 6, we report the win/tie/lose ratio (i.e., “win” means captions generated by MLLMs with CRO is better) and an additional measure of consistency, calculated by the agreement rate between GPT-4o and human judgments. As can be seen, Qwen2.5-VL-3B trained with CRO demonstrates superior caption quality under human evaluation. This aligns with the result in Figure 7, which is further supported by the high consistency of 85%. This supports the rationale of using GPT-4o as a judge for evaluating caption quality.
Table 6: Human evaluation on pairwise comparison of the caption quality.
Win Tie Lose Consistency
60% 29% 11% 85%
C.3 Training Dynamics of CRO
We show the average reward scores and caption lengths over training steps in Figure 13. We observe the following:
  • Rewards increased as training progresses. This confirms the effectiveness of CRO as it allows the MLLM to generate captions that lead to higher reasoning accuracy.
  • Caption lengths grow as training progresses. An explanation for this phenomenon is that the MLLM learns to generate more comprehensive captions during training, which is reflected by longer lengths. This is also confirmed in the Appendix D, where we visualize the captions.
(a) Qwen2.5-VL-3B
(b) Qwen2.5-VL-7B
(c) Qwen2.5-VL-32B
Figure 13: Training dynamics of the CRO process: Reward and Length Over Steps for Qwen2.5-VL-3B, Qwen2.5-VL-7B and Qwen2.5-VL-32B
Appendix D Case Study
Caption qualities.
We conduct a case study on the generated query-conditioned captions. Specifically, for a multi-modal reasoning question and image, we investigate the quality of the generated captions. For MLLMs, we consider Qwen2.5-VL series (3B/32B) both with and without CRO. We visualize the question, image and captions in Tables LABEL:tbl:case_3b_1- LABEL:tbl:case_3b_7 for Qwen2.5-VL-3B and Tables LABEL:tbl:case_32b_1- LABEL:tbl:case_32b_5 for Qwen2.5-VL-32B.
Comparing the captions generated by MLLMs with and without CRO, we discover the following:
  • CRO leads to more visual details. We highlight these visual details in red in the table. Notably, these details are important clues required to correctly solve the question. This shows that CRO is effective in improving the quality (especially comprehensiveness) of the query-conditioned captions.
  • CRO leads to captions with more organized and hierarchical structures. For example, in Table LABEL:tbl:case_3b_6, the MLLM with CRO describes the images at three levels, i.e., Tropic level, Terrestrial food chain and aquatic food chain. This allows the reasoner to quickly locate important information in the captions. However, the original MLLM uses a sequence of sentences that are less clear.
  • Large-sized MLLMs generate more comprehensive captions. We found the captions generated by Qwen2.5-VL-32B are significantly longer than those generated by the 3B model. This is because larger MLLMs have better reasoning abilities that allow it to describe the image from multiple perspectives and in a more logically coherent way. This leads to longer captions.
Reasoning accuracies.
In Table LABEL:tbl:r_case_3b_1, we provide a complete comparison of captions (generated by Qwen2.5-VL-3B with and without CRO) and the resulting reasoning results (produced by R1-7B). Similarly to the case study on caption quality, MLLMs trained with CRO generate captions that capture more details. This is critical to the correctness of the reasoning process. As can be seen, the reasoner that receives captions with CRO arrives at the correct answer after several rounds of thinking and reflection. However, the reasoner that accepts the under-optimized caption experiences multiple contradictions and confusion (highlighted in brown), which leads to responses that exceeds the maximum context length and finally fails this problem.
Table 7: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
If a driver travels an average of 4,250 miles per month driving only along motorways in an Xtam car, what is the predicted annual consumption of fuel (in gallons)? Select from A, B, C, D and E. (A)Cannot say (B) 1,500 (C) 125 (D) 150 (E) 1,250
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
The image is a bar graph titled “Fuel Consumption (miles to the gallon)” that compares the fuel efficiency of five different cars, labeled as Taber, Ursa, Velvo, Tink, and Xtam. The graph is divided into two categories: “City driving” and “Motorway driving.” For each car, there are two bars:
- The light blue bar represents fuel consumption for city driving.
- The dark blue bar represents fuel consumption for motorway driving. The data for each car is as follows: - Taber: City driving: 48 miles per gallon Motorway driving: 43 miles per gallon - Ursa: City driving: 39 miles per gallon Motorway driving: 43 miles per gallon - Velvo: City driving: 35 miles per gallon Motorway driving: 33 miles per gallon - Tink: City driving: 33 miles per gallon Motorway driving: 28 miles per gallon - Xtam: City driving: 33 miles per gallon Motorway driving: 33 miles per gallon Below the bar graph, there is a table that provides additional information about each car:
- Max. Speed (mph): Taber (65), Ursa (60), Velvo (125), Tink (95), Xtam (110)
- Cost to Purchase (£): Taber (12,500), Ursa (15,250), Velvo (37,500), Tink (55,250), Xtam (62,500) The image does not contain any information related to the question about fuel consumption for a specific car model.
Query-conditioned Captions (Qwen2.5-VL-3B):
The image is a bar graph titled “Fuel Consumption (miles to the gallon)” that compares the fuel efficiency of five different cars, Taber, Ursa, Velvo, Tink, and Xtam, under two driving conditions: city driving and motorway driving. The graph uses two colors to differentiate between these conditions: light blue for city driving and dark blue for motorway driving. Below the graph, there is a table that provides additional information about each car, including:
- Max. Speed (mph): The maximum speed each car can achieve.
- Cost to Purchase (£): The price of each car in pounds. The table entries are as follows:
- Taber:
Max. Speed: 65 mph
Cost to Purchase: £12,500
- Ursa:
Max. Speed: 60 mph
Cost to Purchase: £15,250
- Velvo:
Max. Speed: 125 mph
Cost to Purchase: £37,500
- Tink:
Max. Speed: 95 mph
Cost to Purchase: £55,250
- Xtam:
Max. Speed: 110 mph
Cost to Purchase: £62,500 The graph shows that the Taber and Ursa cars have higher fuel efficiency in city driving compared to motorway driving, while the Velvo, Tink, and Xtam cars have higher fuel efficiency in motorway driving compared to city driving.
Table 8: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
According to the question shown in the image, please directly answer the question and provide the correct option letter, e.g., A, B, C, D.
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
The image is a geometric diagram with the following elements:
- Two parallel lines labeled $AB$ and $CD$.
- A line segment $EF$. - A point $G$ where line segment $EG$ intersects line segment $CD$. - An angle labeled $50^{\circ}$ at point $F$.
- An angle labeled $2$ at point $G$. - A question asking about the measure of angle $2$, with choices: A: $50^{\circ}$ B: $60^{\circ}$ C: $65^{\circ}$ D: $90^{\circ}$ The diagram shows a triangle $\triangle EFG$ with angle $1$ at point $F$ and angle $2$ at point $G$. The angle at point $F$ is given as $50^{\circ}$.
The question asks for the measure of angle $2$, given that $AB\parallel CD$ and $EG$ bisects angle $\angle BEF$.
Query-conditioned Captions (Qwen2.5-VL-3B):
The image shows a geometric figure with lines AB and CD parallel to each other. Line EG bisects angle BEF. Angle 1 is marked as 50 degrees.
Table 9: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
Input binary number B = 11000, find the corresponding decimal value by the algorithm shown in image.
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
The image is a flowchart titled "Start Binary To Decimal." It outlines a process to convert a binary number to its decimal equivalent. The flowchart starts with the instruction "Get Binary Number Call it B," which is followed by initializing variables D and Power to 0. The flowchart then checks if B is greater than 0. If B is greater than 0, it checks if B modulo 10 raised to the power of (Power + 1) is equal to 0. If this condition is false, it updates D by adding 2 raised to the power of Power, and then updates B by subtracting 10 raised to the power of Power. If the condition is true, it increments Power. This process continues until B is no longer greater than 0. Finally, it states "Say The Answer is D" and "Stop."
The image does not contain any visual elements or additional information beyond the flowchart itself.
Query-conditioned Captions (Qwen2.5-VL-3B):
Start Binary To Decimal
Get Binary Number Call it B
$D=0$
$\text{Power}=0$
$B>0$
$B\%{10}^{(}\text{Power}+1)=0$
$D=D+2^{\text{Power}}$
$B=B-10^{\text{Power}}$
Yes
Power ++
Table 10: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
The pie chart below represents the number of each grade at a United States elementary school. What is the percentage of Grade 2? The answer should be a percentage between 0 and 100.
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
The image is a pie chart titled "Number of Students in Each Grade." The chart is divided into five segments, each representing a different grade level and the number of students in that grade. The grades and their corresponding numbers are as follows:
- Grade 1: 79 students
- Grade 2: 96 students
- Grade 3: 73 students
- Grade 4: 58 students
- Grade 5: 91 students
The pie chart visually represents the distribution of students across these grades. The segments are color-coded for each grade, with Grade 1 in red, Grade 2 in blue, Grade 3 in green, Grade 4 in orange, and Grade 5 in purple. The chart provides a clear visual representation of the number of students in each grade at the elementary school.
Query-conditioned Captions (Qwen2.5-VL-3B):
The pie chart below represents the number of each grade at a United States elementary school. What is the percentage of Grade 2? The answer should be a percentage between 0 and 100.
Table 11: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
Is the left graph isomorphic with the right graph? choice: (A) Yes (B) No
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
The image consists of two graphs. The left graph is a connected graph with the following vertices and edges: A, B, C, D, E, F, G, and H. The edges are as follows: A is connected to H, A is connected to F, H is connected to E, B is connected to F, F is connected to E, E is connected to G, F is connected to G, and E is connected to D.
The right graph is a simple cycle graph with 8 vertices, labeled as 1, 2, 3, 4, 5, 6, 7 and 8 and edges connecting each vertex to the next in a circular manner.
Based on the description, the left graph is not isomorphic to the right graph. The left graph has a more complex structure with multiple connections and a central node (E), while the right graph is a simple cycle with no such central node and a different number of vertices.
Query-conditioned Captions (Qwen2.5-VL-3B):
The left graph is a graph with 8 vertices and 10 edges. The right graph is a graph with 8 vertices and 8 edges. The left graph is not isomorphic to the right graph.
Table 12: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question: Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end. Question: What happens to fish if pelicans increase?Choices: (A) decrease (B) nothing (C) increase (D) none of the above
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
The image depicts a food chain diagram showing the trophic levels and different types of consumers in both terrestrial and aquatic environments. Here is a detailed description of the image: 1. Trophic Level: - Primary producer (autotroph): This level includes plants and phytoplankton, which are the base of the food chain. - Primary consumer (herbivore): This level includes caterpillars and zooplankton, which feed on the primary producers. - Secondary consumer (carnivore): This level includes lizards and fish, which feed on the primary consumers. - Tertiary consumer (secondary carnivore): This level includes snakes and pelicans, which feed on the secondary consumers. 2. Terrestrial food chain:
- Primary producer (autotroph): Plant
- Primary consumer (herbivore): Caterpillar
- Secondary consumer (carnivore): Lizard
- Tertiary consumer (secondary carnivore): Snake 3. Aquatic food chain:
- Primary producer (autotroph): Phytoplankton
- Primary consumer (herbivore): Zooplankton
- Secondary consumer (carnivore): Fish
- Tertiary consumer (secondary carnivore): Pelican The image also shows arrows indicating the direction of energy flow from one trophic level to the next, emphasizing the flow of energy through different ecosystems.
Query-conditioned Captions (Qwen2.5-VL-3B): The image shows a food chain diagram with two types of food chains: terrestrial and aquatic. The terrestrial food chain includes a caterpillar, lizard, fish, and snake, with the caterpillar being the primary producer and the snake being the tertiary consumer. The aquatic food chain includes zooplankton, phytoplankton, fish, and a pelican, with the zooplankton being the primary producer and the pelican being the tertiary consumer.
Table 13: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end. Question: If frogs were removed from this environment what animal would potentially see an increase in its population? Choices:(A) crickets (B) deer (C) snakes (D) hawks
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
The image is a food web diagram depicting various organisms and their interactions within an ecosystem. The diagram includes the following elements: Deer (at the bottom left) consuming Trees. Mice (in the center) consuming Grasses. Crickets (in the center right) consuming Grasses. Snakes (at the top left) consuming Mice. Hawks (at the top center) consuming Snakes. Owls (at the top right) consuming Mice. Frogs (in the center right) consuming Crickets. Decomposers (at the bottom center) breaking down organic matter. The arrows indicate the direction of energy flow from one organism to another, showing the relationships between the different species in the ecosystem.
Query-conditioned Captions (Qwen2.5-VL-3B): The image is a food web diagram depicting various organisms and their interactions within an ecosystem. It includes a deer, trees, grasses, mice, crickets, frogs, hawks, and owls. The arrows indicate the direction of energy flow from producers (trees and grasses) to consumers (deer, mice, crickets, frogs, hawks, and owls), and decomposers (decomposers). The diagram shows how energy moves through the ecosystem, with each organism feeding on others, and how the removal of frogs could potentially lead to an increase in the population of crickets, as they are a primary food source for frogs.
Table 14: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end. Question: Which option is correct? Choices: (A) A (B) B (C) C
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
The image presents a logic puzzle involving patterns and shapes. Here is a detailed description of the image: Left Side: Grid with Shapes The grid is a 3x3 matrix. Each cell contains a shape with a specific pattern or design. The shapes include circles, diamonds, and triangles, each with different internal designs (e.g., stars, squares, etc.). Row 1: 1. Cell 1: A circle with a star inside. 2. Cell 2: A diamond with a star inside. 3. Cell 3: A triangle with a star inside. Row 2: 1. Cell 1: Empty. 2. Cell 2: A circle with a square inside. 3. Cell 3: A diamond with a square inside. Row 3: 1. Cell 1: Empty. 2. Cell 2: A triangle with a downward arrow inside. 3. Cell 3: Empty. Right Side: Multiple Choice Options There are three options labeled A, B, and C. Each option shows a sequence of three shapes. The task is to determine the correct pattern that fits the logic of the grid. Option A: 1. A diamond with a downward arrow inside. 2. A triangle with a square inside. 3. A circle with a star inside. Option B: 1. A diamond with a downward arrow inside. 2. A triangle with a square inside. 3. A circle with a downward arrow inside. Option C:
1. A diamond with a downward arrow inside. 2. A triangle with a square inside. 3. A diamond with a downward arrow inside. Analysis of the Grid Patterns Column-wise Observation: Column 1: Row 1: Circle with a star. Row 2: Empty. Row 3: Empty. Column 2: Row 1: Diamond with a star. Row 2: Circle with a square. Row 3: Triangle with a downward arrow. Column 3: Row 1: Triangle with a star. Row 2: Diamond with a square. Row 3: Empty (to be determined). Row-wise Observation: Row 1: Circle (star), Diamond (star), Triangle (star) Row 2: Empty, Circle (square), Diamond (square) Row 3: Empty, Triangle (downward arrow), ? (to be determined) Shape Progression: The shapes in each row and column follow a specific order: Circle $\rightarrow$ Diamond $\rightarrow$ Triangle. The internal designs also follow a pattern: Star $\rightarrow$ Square $\rightarrow$ Downward Arrow.
Query-conditioned Captions (Qwen2.5-VL-32B):
The image consists of two main sections: Left Section: A 3x3 grid with various shapes and patterns. The grid contains: A circle with a star inside in the top-left cell. A diamond with a star inside in the top-middle cell. A triangle with a star inside in the top-right cell. A circle with a square inside in the middle-left cell. A diamond with a square inside in the middle-middle cell. A triangle with a square inside in the middle-right cell. A triangle with a downward-pointing arrow inside in the bottom-left cell. The bottom-middle and bottom-right cells are empty.
Right Section: Three options labeled A, B, and C. Each option contains three shapes: Option A: A diamond with a downward-pointing arrow inside, a triangle with a square inside, and a circle with a star inside. Option B: A circle with a downward-pointing arrow inside, a triangle with a square inside, and another circle with a downward-pointing arrow inside. Option C: A circle with a downward-pointing arrow inside, a triangle with a square inside, and a diamond with a downward-pointing arrow inside. The task appears to involve identifying a pattern or sequence based on the shapes and their internal symbols in the grid and selecting the correct option from the choices provided.
Table 15: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
Hint: Please answer the question requiring an integer answer and provide the final value, e.g., 1, 2, 3, at the end. Question: How many models in the figure achieve an Acc score greater than 60?
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
The image is a scatter plot titled "Model Performance," which compares the performance of various large language models on a specific task, likely related to the GSM8K dataset. The plot uses different colors and line styles to distinguish between models and their training settings. Here is a detailed description of the image: Title and Labels: Title: "Model Performance" X-axis: Labeled "Pretrain Loss," ranging from approximately 1.2 to 2.0. Y-axis: Labeled "ACC" (presumably short for Accuracy), ranging from 0 to 80. Legend: The legend on the left side lists various models and their corresponding colors and markers: GPT3-3B (pink), GPT3-6B (orange), GPT3-12B (yellow), GPT3-175B (dark green) LLaMA-7B (light green), LLaMA2-7B (teal), LLaMA-13B (dark teal), LLaMA2-13B (light blue) LLaMA-33B (blue), LLaMA-65B (purple), LLaMA2-70B (pink), GPT-4 (dark pink) Line Styles: Blue Lines: Represent the SFT (Supervised Fine-Tuning) setting. Red Lines: Represent the ICL (In-Context Learning) setting. Dotted lines connect points for each model, showing trends in performance. Data Points: Each model is represented by a series of points on the plot, with the position determined by: X-axis (Pretrain Loss): Lower values indicate better pretraining. Y-axis (ACC): Higher values indicate better accuracy. Key Observations: 1. GPT-4 Performance: GPT-4 (dark pink) achieves the highest accuracy (ACC) score, well above 60, and has a low pretraining loss. It is plotted as a single point at the top-right of the graph. 2. LLaMA and LLaMA2 Models: Various LLaMA and LLaMA2 models (e.g., LLaMA-7B, LLaMA2-7B, LLaMA-13B, etc.) are scattered across the plot. Larger models (e.g., LLaMA2-70B) tend to have higher accuracy scores. 3. GPT3 Models: GPT3 models (GPT3-3B, GPT3-6B, GPT3-12B, GPT3-175B) are plotted with lower accuracy scores compared to LLaMA and LLaMA2 models. GPT3-175B performs better than smaller GPT3 models but still lags behind the larger LLaMA models. 4. Accuracy Threshold (ACC > 60): Models achieving an ACC score greater than 60 are primarily from the LLaMA2 series and GPT-4. Notable models include: LLaMA2-70B LLaMA2-65B LLaMA2-33B GPT-4 Text Caption: The caption below the figure provides additional context: Figure 2: Describes the performance of SFT (blue lines) and ICL (red lines) settings on GSM8K. Notes that GPT-4 uses some part of the GSM8K data in pretraining and suggests its performance lies between SFT and ICL.
Query-conditioned Captions (Qwen2.5-VL-32B):
The image is a scatter plot titled "Model Performance." It compares the performance of various large language models on a task, likely related to the GSM8K dataset, based on two metrics: "ACC" (accuracy) on the y-axis and "Pretrain Loss" on the x-axis. The models are represented by different colored dots, and the legend on the left identifies each model by name and color. The models include various versions of GPT-3 (3B, 6B, 12B, 175B), LLaMA (7B, 13B, 33B, 65B, 70B), LLaMA2 (7B, 13B, 33B, 65B, 70B), and GPT-4. Key Features: 1. Axes: Y-axis (ACC): Ranges from 0 to 80, representing accuracy scores. X-axis (Pretrain Loss): Ranges from 1.2 to 2.0, representing pretraining loss values. 2. Models: Each model is represented by a colored dot, with the model name labeled near the dot. Models include the GPT-3 series, LLaMA series, LLaMA2 series, and GPT-4. 3. Trends: There are two sets of lines: Blue lines (SFT settings): Represent the performance of models under supervised fine-tuning (SFT). Red lines (ICL settings): Represent the performance of models under in-context learning (ICL). GPT-4 is shown as a single point, suggesting its unique position in the performance landscape. 4. Legend: The legend on the left provides a color-coded list of models, including: GPT-3 variants: 3B, 6B, 12B, 175B LLaMA variants: 7B, 13B, 33B, 65B, 70B LLaMA2 variants: 7B, 13B, 33B, 65B, 70B GPT-4 5. Caption: The caption explains that the figure shows the performance of SFT (blue lines) and ICL (red lines) settings on GSM8K. It notes that GPT-4 uses some part of the GSM8K data in pretraining and suggests its performance lies between SFT and ICL. Question: The question asks how many models in the figure achieve an ACC score greater than 60. The task is to count the number of models that meet this criterion based on the scatter plot. However, the instruction is to describe the image without answering the question.
Table 16: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end. Question: Does South Carolina have the highest value in the South ? Choices: (A) Yes (B) No
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
The image is a choropleth map of the United States, depicting the distribution of health insurance coverage for nonelderly individuals (ages 0–64) through employer-sponsored plans in 2013. Here is a detailed description of the image: Title and Subtitle: Title: “Health Insurance Coverage of Nonelderly 0–64: Employer, 2013” This indicates that the map shows the number of nonelderly individuals (aged 0–64) who had employer-sponsored health insurance in 2013. Color Coding: The map uses a gradient of blue shades to represent different ranges of employer-sponsored health insurance coverage. The legend at the bottom explains the color coding: Lightest Blue: 281,900 – 2,024,500 Light Blue: 2,098,300 – 4,739,300 Medium Blue: 4,757,100 – 6,467,400 Darkest Blue: 7,133,900 – 16,665,500 State Colors: Darkest Blue (Highest Coverage): States in this category include California (CA), New York (NY), Texas (TX), and Florida (FL). These states have the highest number of nonelderly individuals with employer-sponsored health insurance. Medium Blue: States like Illinois (IL), Ohio (OH), Pennsylvania (PA), and Michigan (MI) fall into this category. Light Blue: States such as Colorado (CO), Arizona (AZ), and Tennessee (TN) are in this range. Lightest Blue (Lowest Coverage): States with the lowest coverage include Alaska (AK), Hawaii (HI), Wyoming (WY), and several others in the Midwest and Northeast.
South Region Analysis: Highest Value in the South: The darkest blue states in the South are Texas (TX) and Florida (FL), indicating they have the highest employer-sponsored health insurance coverage in the region. South Carolina (SC): South Carolina is colored in a light blue shade, indicating it falls into the range of 2,098,300 – 4,739,300. This is not the highest value in the South. Other Observations: Northeast: New York (NY) and Pennsylvania (PA) have high coverage, with NY in the darkest blue category. Midwest: States like Illinois (IL) and Ohio (OH) have medium to high coverage. West: California (CA) stands out with the darkest blue, indicating the highest coverage in the West. Source: The map is sourced from the Kaiser Family Foundation’s State Health Facts, as noted at the bottom of the image. Question Analysis: The question asks whether South Carolina has the highest value in the South. Based on the map: South Carolina is in a light blue category, not the darkest blue. The highest values in the South are represented by Texas and Florida, which are in the darkest blue category.
Query-conditioned Captions (Qwen2.5-VL-32B):
The image is a choropleth map of the United States, illustrating the distribution of health insurance coverage for nonelderly individuals aged 0–64 through employer-sponsored plans in 2013. The map uses varying shades of blue to represent different ranges of coverage numbers across the states. The legend at the bottom indicates four categories of coverage: Lightest Blue: 281,900 – 2,024,500 Light Blue: 2,098,300 – 4,739,300 Medium Blue: 4,757,100 – 6,467,400 Darkest Blue: 7,133,900 – 16,665,500 Key observations from the map: States like California (CA), Texas (TX), and Florida (FL) are shaded in the darkest blue, indicating the highest range of employer-sponsored health insurance coverage. Many states in the Midwest and West are shaded in lighter blues, indicating lower coverage numbers. The source of the data is cited as the Kaiser Family Foundation’s State Health Facts. The question asks whether South Carolina (SC) has the highest value in the South. The map shows South Carolina in a light blue shade, indicating it falls into the second-lowest category of coverage.
Table 17: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
<image 1>A battery, an ammeter, three resistors, and a switch are connected to form the simple circuit shown above. When the switch is closed what would happen to the potential difference across the 15 ohm resistor?
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
The image depicts an electrical circuit involving a battery, an ammeter, three resistors, and a switch. Here is a detailed description of the components and their arrangement: Components in the Circuit: 1. Battery: Located on the left side of the circuit. Serves as the power source, providing the electromotive force (EMF) to the circuit. 2. Ammeter (A): Placed immediately after the battery. Used to measure the current flowing through the circuit. It is connected in series with the rest of the circuit. 3. Resistors: There are three resistors with the following resistances: A 20 $\Omega$ resistor A 15 $\Omega$ resistor A 30 $\Omega$ resistor The resistors are connected in a combination of series and parallel configurations. 4. Switch (S): Located on the right side of the circuit. Controls the flow of current in the circuit. When closed, the circuit is complete, and current can flow. Circuit Configuration: The 20 $\Omega$ resistor is connected in series with the rest of the circuit. The 15 $\Omega$ and 30 $\Omega$ resistors are connected in parallel. The parallel combination is then connected in series with the 20 $\Omega$ resistor. The ammeter is in series with all the resistors, measuring the total current. The switch is placed after the parallel combination, completing the circuit when closed.
Path of Current: 1. Current flows from the positive terminal of the battery. 2. It passes through the ammeter. 3. It then encounters the 20 $\Omega$ resistor. 4. After the 20 $\Omega$ resistor, the current splits: One part flows through the 15 $\Omega$ resistor. The other part flows through the 30 $\Omega$ resistor. 5. The currents recombine after the parallel branch. 6. The combined current continues through the rest of the circuit and returns to the battery. Analysis of Potential Difference: When the switch is closed, current flows through the circuit and potential differences develop across each resistor. The 15 $\Omega$ and 30 $\Omega$ resistors are in parallel. In a parallel configuration, the potential difference across each branch is the same. Therefore, the potential difference across the 15 $\Omega$ resistor is equal to that across the 30 $\Omega$ resistor.
Query-conditioned Captions (Qwen2.5-VL-32B):
The image depicts a simple electrical circuit consisting of the following components: Battery: A power source labeled as "Battery" is shown on the left side of the circuit. Ammeter (A): A device used to measure the current in the circuit is connected in series with the battery. Resistors: Three resistors are present in the circuit: A 20 $\Omega$ resistor is connected in series with the ammeter. A 15 $\Omega$ resistor is connected in parallel with a 30 $\Omega$ resistor. Switch (S): A switch is placed in the circuit, which can be opened or closed to control the flow of current. Wires: The components are connected by wires to form a complete circuit. Circuit Configuration: The 20 $\Omega$ resistor is in series with the ammeter and the battery. The 15 $\Omega$ resistor and the 30 $\Omega$ resistor are connected in parallel with each other. The switch is placed in the branch that includes the parallel combination of the 15 $\Omega$ and 30 $\Omega$ resistors. Key Observations: When the switch is closed, the parallel combination of the 15 $\Omega$ and 30 $\Omega$ resistors will be active, allowing current to flow through both resistors. The ammeter measures the total current flowing through the circuit. This description provides a detailed overview of the circuit’s components and their arrangement.
Table 18: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
Image:
Question:
Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end.Question: Misty the cat has five kittens: two of them are striped, one spotty, the rest of them are absolutely white. In which picture can we see the kittens of Misty, knowing that the ears of one of them are of different colour? Choices: (A) A (B) B (C) C (D) D (E) E
Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
The image presents a multiple-choice question involving illustrations of cats. Here is a detailed description of the image: Overall Structure: The image contains five labeled sections, each representing a different option: A, B, C, D, and E. Each section features a group of cartoon-style cat illustrations. The question describes a specific scenario about a cat named Misty who has five kittens with particular characteristics. The task is to identify which section correctly represents Misty’s kittens based on the given clues. Clues from the Question: 1. Misty has five kittens. 2. Two of the kittens are striped. 3. One kitten is spotty. 4. The rest (two kittens) are absolutely white. 5. One of the kittens has ears of a different color. Analysis of Each Section: Section A: 1 striped, 1 spotty, 2 white kittens. No kitten has ears of a different color. Section B: 2 striped, 1 spotty, 2 white kittens. No kitten with differently colored ears. Section C: 2 striped, 1 spotty, 2 white, 1 gray kittens. One kitten has ears of a different color.
Section D: 2 striped, 1 spotty, 2 white kittens. One kitten has ears of a different color. Section E: 1 striped, 3 spotty, 1 white kittens. One kitten has ears of a different color.
Query-conditioned Captions (Qwen2.5-VL-32B):
[…]
Table 19: Reasoning Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red. Reasoning process that show confusing and contradiction are highlighted in brown.
Image:
Question:
O is the circumcircle of $\triangle ABC$, with $AB=BC=16$. The arc $AB$ is folded down along the chord $AB$ to intersect $BC$ at point $D$. If point $D$ is the midpoint of $BC$, what is the length of $AC$?
Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
The image is a geometric diagram of a triangle $\triangle ABC$ inscribed in a circle with center $O$. The diagram shows the following elements:
- Points $A$, $B$, and $C$ are the vertices of the triangle.
- Point $O$ is the center of the circumcircle of $\triangle ABC$.
- Line segment $AB$ is a chord of the circle.
- Line segment $BC$ is another chord of the circle.
- Point $D$ is a point on the circle such that the arc $AB$ is folded down along the chord $AB$ to intersect $BC$ at $D$.
- The diagram includes a dashed line segment $AD$ and a dashed line segment $BD$.
The image does not provide any numerical values or specific measurements, and it does not contain any text or additional information. The focus is purely on the geometric relationships and the diagram itself.
Query-conditioned Captions (Qwen2.5-VL-3B):
The image shows a circle with center $O$ and points $A$, $B$, and $C$ on its circumference. The line segment $AB$ is a diameter of the circle, and the arc $AB$ is folded down along the chord $AB$ to intersect $BC$ at point $D$. Point $D$ is the midpoint of $BC$.
Reasoning Result of R1-7B (captions from Qwen2.5-VL-3B w/ CRO):
Alright, let me try to figure out this geometry problem step by step. So, we have triangle $ABC$ inscribed in a circle with center $O$. $AB$ and $BC$ are both 16 units long, making it an isosceles triangle with $AB=BC$. The diagram shows that when arc $AB$ is folded along chord $AB$, it intersects $BC$ at point $D$, which is the midpoint of $BC$. So, $BD=DC=8$ units.
First, I need to visualize this. Since $ABC$ is isosceles with $AB=BC$, the base is $AC$, and the two equal sides are $AB$ and $BC$. The circle is the circumcircle, so all three vertices lie on the circumference.
When the arc $AB$ is folded along $AB$, point $D$ is where the fold intersects $BC$. Since $D$ is the midpoint, $BD=DC=8$. Folding along $AB$ suggests that point $D$ is the reflection of point $C$ over the line $AB$. But since $D$ lies on $BC$, maybe there’s some symmetry here.
I think I should consider the properties of the circumcircle. In a triangle inscribed in a circle, the perpendicular bisectors of the sides meet at the center $O$. Since $AB=BC$, the triangle is isosceles, so the median from $B$ to $AC$ should also be the altitude and the angle bisector.
Wait, but we’re dealing with a fold along $AB$. When you fold along $AB$, point $C$ would map to some point on $BC$. Since $D$ is the midpoint, maybe there’s a reflection involved here. Let me try to see if triangle $ABD$ is congruent or similar to another triangle.
Alternatively, maybe using the Law of Cosines would help. Let’s denote angle at $B$ as angle $ABC$. Since $AB=BC=16$, the triangle is isosceles, so angles at $A$ and $C$ are equal. But when we fold arc $AB$ over $AB$, point $D$ is created on $BC$. Since $D$ is the midpoint, perhaps there’s a relationship between the angles or the lengths.
Wait, another approach: when you fold along $AB$, point $C$ maps to $D$. So, $AD$ should equal $AC$ because reflection preserves distances. So, $AD=AC$. Also, since $D$ is the midpoint of $BC$, $BD=DC=8$. So, in triangle $ABD$, we have $AB=16$, $BD=8$, and $AD=AC$.
So, triangle $ABD$ has sides $AB=16$, $BD=8$, and $AD=AC=x$. Maybe I can apply the Law of Cosines here. Let’s denote $AC$ as $x$. Then, $AD=x$ as well.
In triangle $ABD$: $AB^{2}+BD^{2}-2\cdot AB\cdot BD\cdot\cos(\angle ABD)=AD^{2}$ Which is: $16^{2}+8^{2}-2\cdot 16\cdot 8\cdot\cos(\angle ABD)=x^{2}$ $256+64-256\cdot\cos(\angle ABD)=x^{2}$ $320-256\cdot\cos(\angle ABD)=x^{2}$
But I don’t know angle $ABD$. Maybe I can relate it to the original triangle $ABC$. Since $ABC$ is isosceles with $AB=BC=16$, angle $ABC$ is equal to angle $BAC$ and angle $BCA$? Wait, no, in an isosceles triangle, the angles opposite the equal sides are equal. So, since $AB=BC$, the angles opposite them, which would be angle $ACB$ and angle $BAC$, are equal.
Wait, no, in triangle $ABC$, $AB=BC=16$, so the sides opposite angles at $C$ and $A$ respectively. So, angle at $C$ (angle $ACB$) equals angle at $A$ (angle $BAC$). Let’s denote angle $BAC$ as $\theta$, so angle $ACB$ is also $\theta$, and angle $ABC$ is $180-2\theta$ degrees.
When we fold along $AB$, point $C$ maps to $D$ on $BC$. So, the reflection over $AB$ takes point $C$ to $D$. Therefore, $AD=AC$ and angle $BAD=$ angle $BAC=\theta$.
Wait, but $AD$ is equal to $AC$ because reflection preserves length. So, triangle $ABD$ has sides $AB=16$, $BD=8$, and $AD=AC=x$. Maybe I can use the Law of Cosines here as well.
Alternatively, maybe using the Law of Sines in triangle $ABC$. Since $ABC$ is isosceles with $AB=BC=16$, the circumradius can be found using the formula: $R=\frac{a}{2\cdot\sin(A)}$ Where $a$ is the length of a side, and $A$ is the opposite angle.
Wait, but I’m not sure if that’s the right approach. Let me think again.
Since $ABC$ is isosceles with $AB=BC=16$, then angle at $A$ equals angle at $C$. Let me denote angle at $B$ as $\phi$. So, in triangle $ABC$, angles are $\phi$ at $B$, and $(180-\phi)/2$ at $A$ and $C$ each.
Now, when we fold along $AB$, point $C$ is reflected over $AB$ to point $D$ on $BC$. So, $AD=AC$, and $BD=8$, $DC=8$.
So, in triangle $ABD$, $AB=16$, $BD=8$, $AD=x$.
So, using the Law of Cosines on triangle $ABD$: $AB^{2}+BD^{2}-2\cdot AB\cdot BD\cdot\cos(\angle ABD)=AD^{2}$ $16^{2}+8^{2}-2\cdot 16\cdot 8\cdot\cos(\angle ABD)=x^{2}$ $256+64-256\cdot\cos(\angle ABD)=x^{2}$ $320-256\cdot\cos(\angle ABD)=x^{2}$
But angle $ABD$ is the same as angle $ABC$, which is $\phi$. So, $\cos(\angle ABD)=\cos(\phi)$.
From triangle $ABC$, using the Law of Cosines: $AC^{2}=AB^{2}+BC^{2}-2\cdot AB\cdot BC\cdot\cos(\phi)$ $x^{2}=16^{2}+16^{2}-2\cdot 16\cdot 16\cdot\cos(\phi)$ $x^{2}=512-512\cdot\cos(\phi)$
But from triangle $ABD$, we have $x^{2}=320-256\cdot\cos(\phi)$
So, setting them equal: $512-512\cdot\cos(\phi)=320-256\cdot\cos(\phi)$ $512-320=512\cdot\cos(\phi)-256\cdot\cos(\phi)$ $192=256\cdot\cos(\phi)$ $\cos(\phi)=\frac{192}{256}=\frac{3}{4}$
So, angle $\phi$ has cosine $\frac{3}{4}$. Now, substituting back into the equation for $x^{2}$ from triangle $ABC$: $x^{2}=512-512\cdot\left(\frac{3}{4}\right)$ $x^{2}=512-384=128$ $x=\sqrt{128}=8\sqrt{2}$
Wait, but this contradicts the initial thought that $AC$ is $16\sqrt{2}$. Hmm, maybe I made a mistake in the reflection part. Let me check.
Wait, when we fold along $AB$, point $C$ maps to $D$ on $BC$. So, $AD$ should equal $AC$ because reflection preserves distance. So, $AD=AC=x$.
But in triangle $ABD$, we have $AB=16$, $BD=8$, $AD=x$.
Using Law of Cosines: $16^{2}+8^{2}-2\cdot 16\cdot 8\cdot\cos(\angle ABD)=x^{2}$ $256+64-256\cdot\cos(\angle ABD)=x^{2}$ $320-256\cdot\cos(\angle ABD)=x^{2}$
But angle $ABD$ is the same as angle $ABC$, which is $\phi$. So, $\cos(\angle ABD)=\cos(\phi)=\frac{3}{4}$. Thus, $x^{2}=320-256\cdot\left(\frac{3}{4}\right)=320-192=128$ $x=\sqrt{128}=8\sqrt{2}$
Wait, so that would make $AC=8\sqrt{2}$, not $16\sqrt{2}$ as the initial model response suggested. Hmm, so maybe the model response was incorrect.
Alternatively, perhaps I made a mistake in assuming angle $ABD$ is equal to angle $ABC$. Let me think again.
In triangle $ABC$, angle $ABC$ is at point $B$, between $AB$ and $BC$. When we fold along $AB$, the angle at $B$ remains the same, but the direction changes. So, the angle $ABD$ in triangle $ABD$ is actually angle between $AB$ and $BD$, which is different from angle $ABC$.
Wait, no, $BD$ is part of $BC$, so angle $ABD$ is actually the same as angle $ABC$ because $BD$ is along $BC$. So, maybe my initial assumption was correct.
Wait, but then according to this calculation, $AC=8\sqrt{2}$, which is about 11.31, but given that $AB$ and $BC$ are 16, that seems plausible.
Alternatively, maybe I should consider that when folding, the angle at $D$ is a right angle. So, maybe triangle $ABD$ is a right triangle.
Wait, in the model response, they concluded that angle $BAC$ is 90 degrees, which might not be correct. Let me see.
If angle $BAC$ is 90 degrees, then triangle $ABC$ would be a right-angled isosceles triangle, making $AC=AB\cdot\sqrt{2}=16\sqrt{2}$. But according to my calculation, it’s $8\sqrt{2}$. So, maybe the model response was wrong.
Wait, perhaps the mistake is in assuming that $AD=AC$. When you fold along $AB$, point $C$ maps to $D$, so the distance from $A$ to $D$ should equal the distance from $A$ to $C$, right? So, $AD=AC$. So, that part seems correct.
Wait, but in the model response, they used Law of Cosines with angle $BAC$ being 90 degrees, which led to $AC=16\sqrt{2}$. But according to my calculation, angle $BAC$ is not necessarily 90 degrees, and it’s actually determined by the folding.
Wait, perhaps I should use the fact that in the folded diagram, $AD=AC$ because of reflection. So, maybe triangle $ABD$ is congruent to triangle $ABC$? But no, because $BD$ is only 8, not 16.
Alternatively, perhaps using coordinate geometry would help. Let me place point $A$ at (0,0), point $B$ at (16,0), since $AB=16$. Since $ABC$ is isosceles with $AB=BC=16$, point $C$ would be somewhere such that $BC=16$. … (rest of the reasoning) Therefore, the correct answer should be $AC=8\sqrt{2}$, not $16\sqrt{2}$ as in the model response.
But wait, maybe I made a mistake in assuming angle $ABD$ is equal to angle $ABC$. Let me think again.
In triangle $ABC$, angle at $B$ is $\phi$, which we found $\cos(\phi)=\frac{3}{4}$. When folding along $AB$, the angle at $B$ remains the same, so angle $ABD$ is indeed $\phi$. Therefore, my calculation should be correct.
Therefore, the model response was incorrect, and the correct answer is $\boxed{AC=8\sqrt{2}}$.
Reasoning Result of R1-7B (captions from Qwen2.5-VL-3B):
Alright, so I’m trying to solve this geometry problem.
Let me read it again carefully. We have triangle ABC with AB = BC = 16.
So, it’s an isosceles triangle with AB and BC equal.
That means angles at A and C are equal, right?
Wait, no, actually, in triangle ABC, if AB = BC, then the base is AC, so the base angles at A and C are equal.
Hmm, but I’m not sure if that’s important right now. The problem mentions that the arc AB is folded down along the chord AB to intersect BC at point D.
And D is the midpoint of BC.
So, BC is 16, so BD and DC are each 8.
Got it. We need to find the length of AC. First, I should probably visualize the problem.
Let me draw a circle with center O, points A, B, and C on the circumference.
AB is a diameter, but wait, hold on.
The caption says the image shows a circle with center O and points A, B, C on the circumference.
The line segment AB is a diameter, and the arc AB is folded down along AB to intersect BC at D, which is the midpoint. Wait, so AB is a diameter, which would mean that angle ACB is a right angle because any triangle inscribed in a semicircle is a right triangle.
But hold on, in the problem statement, it says AB = BC = 16.
If AB is a diameter, then AB would be the longest side, but BC is also 16, so that’s interesting because in a right triangle, the hypotenuse is the longest side.
So if AB is the diameter and a side of the triangle, then AB would be the hypotenuse.
But in the problem, AB = BC, so both would be legs, which can’t be because in a right triangle, the hypotenuse is longer than the legs.
Hmm, maybe AB isn’t the diameter? Wait, the initial caption says AB is a diameter, but the problem statement says AB = BC = 16.
That seems conflicting because if AB is a diameter, it should be longer than BC.
Maybe I misread the problem.
Let me check. Wait, no, the problem says O is the circumcircle of triangle ABC.
So O is the circumradius.
If AB is a diameter, then the circumradius would be half of AB, which would make the circumradius 8.
But if AB is 16, then the circumradius is 8.
However, in that case, triangle ABC would have AB as the diameter, making angle ACB a right angle.
But in the problem, AB = BC = 16, so triangle ABC would have two sides equal to 16 and one side (AC) which we need to find.
But if angle C is 90 degrees, then AC would be the other leg, so AC = BC = 16, making it an isosceles right triangle, so AC would be $16\sqrt{2}$.
But wait, that seems too straightforward, and maybe that’s the answer, but I’m not sure because the folding part might complicate things.
Wait, the problem mentions that the arc AB is folded down along AB to intersect BC at D, which is the midpoint.
So maybe the initial assumption that AB is a diameter isn’t correct because when you fold the arc AB down along AB, it intersects BC at D, which is the midpoint, so maybe AB isn’t a diameter.
Hmm, now I’m confused.
Alternatively, let me try to think differently.
Maybe AB isn’t the diameter, but just a chord.
So the circle has center O, and AB is a chord, not necessarily the diameter.
Then, the arc AB is folded along AB, meaning reflecting the arc AB over the chord AB, which would create a reflection of point C over AB, landing at point D on BC. Since D is the midpoint of BC, which is 16, so BD = DC = 8.
So, when you fold arc AB over AB, the reflection of C over AB is D, which lies on BC.
So, perhaps triangle ABD is congruent to triangle ABC, but I’m not sure.
Alternatively, maybe the reflection of C over AB is D, which is on BC.
So, the distance from C to AB is the same as from D to AB, but since D is on BC, maybe we can use some properties here.
Wait, perhaps using coordinates would help.
Let me try to place the points on a coordinate system.
Let’s place point B at the origin (0,0).
Since AB = 16, let me place point A at (16, 0).
Then, since AB is a chord of the circle, the center O is somewhere in the plane.
Let me denote point C as (x, y).
Since AB = BC = 16, the distance from B to C is 16, so $\sqrt{x^{2}+y^{2}}=16$.
Also, since ABC is a triangle, point C is somewhere not on AB. Now, the arc AB is folded over AB, and it intersects BC at D, which is the midpoint of BC.
So, since B is at (0,0) and C is at (x,y), the midpoint D is at $(\frac{x}{2},\frac{y}{2})$. When we fold the arc AB over AB, the reflection of point C over AB is point D.
So, the reflection of point C over AB is D, which lies on BC.
So, the reflection of point C over AB is point D. So, to find the reflection of point C over AB, let’s find the equation of AB first.
Since A is at (16,0) and B is at (0,0), AB is the x-axis from (0,0) to (16,0).
So, the line AB is y=0. Reflecting a point over the x-axis changes the sign of the y-coordinate.
So, the reflection of point C=(x,y) over AB (y=0) would be (x, -y).
But the reflection point is D, which is $(\frac{x}{2},\frac{y}{2})$.
So, according to the reflection, (x, -y) should equal $(\frac{x}{2},\frac{y}{2})$.
Wait, that can’t be unless x = 0 and y = 0, which would place C at the origin, but then ABC wouldn’t be a triangle.
Hmm, that doesn’t make sense. Maybe my assumption that AB is the x-axis is causing confusion because in reality, AB is just a chord, not necessarily the x-axis.
Alternatively, perhaps I should place AB as a horizontal chord but not necessarily aligned with the x-axis.
Let me try that.
Let me place point A at (a, b) and point B at (c, d), but this might complicate things.
Maybe a better approach is to use coordinate geometry with AB as a chord, not necessarily aligned with any axis.
Alternatively, maybe using vectors or geometric properties would be better.
Let’s think about the reflection.
When you fold the arc AB over AB, point C is reflected over AB to point D on BC.
… omitted When we fold arc AB over AB, the reflection of point C over AB is point D.
So, the reflection of C over AB is D.
Since AB is the x-axis from (0,0) to (a,0), the reflection over AB would invert the y-coordinate.
So, if C is at (x,y), then D would be at (x, -y).
But D is also the midpoint of BC, which is $(\frac{x_{B}+x_{C}}{2},\frac{y_{B}+y_{C}}{2})=(\frac{0+x}{2},\frac{0+y}{2})=(\frac{x}{2},\frac{y}{2})$.
So, setting (x, -y) equal to $(\frac{x}{2},\frac{y}{2})$: x = x/2 => x = 0
-y = y/2 => -y = y/2 => -2y = y => 3y = 0 => y = 0 Again, this leads to y=0, which would place C on AB, making ABC degenerate.
Hmm, clearly, placing AB on the x-axis isn’t working because it forces C to be on AB. Let me try a different approach without coordinate geometry. Let me consider the properties of the reflection.
When we fold arc AB over AB, point C is reflected over AB to point D on BC.
So, AB is the perpendicular bisector of segment CD.
Therefore, AB is perpendicular to CD, and the midpoint of CD lies on AB. Since D is the midpoint of BC, the midpoint M of CD is also the midpoint of BC, which is D itself.
So, D lies on AB and is the midpoint of CD.
Therefore, AB is the perpendicular bisector of CD, so CD is perpendicular to AB, and D is the midpoint of CD and lies on AB. Wait, but D is also the midpoint of BC.
So, if D lies on AB, then BC must intersect AB at D, which is the midpoint of BC.
So, in triangle ABC, D is the midpoint of BC and lies on AB.
Therefore, AB is a median of the triangle, and since AB is also a side, it must be that triangle ABC is such that AB is both a side and a median, which is only possible if ABC is an isosceles triangle with AB = AC.
But the problem states AB = BC = 16, so if AB = AC, then it’s an equilateral triangle, but in that case, all sides would be 16, making AC = 16, but the answer seems to be $16\sqrt{2}$, so maybe that’s not the case. I’m really confused now. Maybe the correct answer is $16\sqrt{2}$, as in the initial tentative solution, but I’m not sure why.
Wait, another idea: since D is the midpoint of BC and lies on AB, and AB is a side, then ABC must be such that AB is a median and an altitude, making it an isosceles triangle with AB = AC.
So, AB = AC = 16. … exceeding maximum context length (8192)
Appendix E Broader Impacts
RACRO improves multi-modal reasoning by aligning perception with reasoning through reinforcement learning. While beneficial for tasks requiring precise vision-language understanding, this alignment could be misused for generating misleading content or surveillance. Caution is advised when applying RACRO in sensitive domains. Mitigation strategies include human oversight, responsible deployment, and transparency in model behavior.
Appendix F Limitation
We evaluate the effectiveness of the proposed RACRO framework exclusively on multi-modal reasoning tasks. However, its potential applicability could extend to other important domains, such as medical visual question answering (VQA) and geospatial reasoning. Exploring these areas represents promising directions for future work.
Appendix G More Discussions
Future works.
This work aims at decoupling multi-modal reasoning with visual perception and LLM reasoning, and the successful visual perception is essential for downstream reasoning. How to further enhance the visual perception capability is a prevailing research direction (e.g., utilizing AIGC-generated data [10, 16, 17, 18, 23, 28, 29, 33, 50] and enhanced visual representations [7, 8, 35, 58])

### v2

@@
-Perceptual Decoupling for Scalable Multi-modal Reasoning via Reward-Optimized Captioning
+Reasoning-Aligned Perception Decoupling for Scalable Multi-modal Reasoning
 Abstract
-Recent advances in slow-thinking language models (e.g., OpenAI-o1 and DeepSeek-R1) have demonstrated remarkable abilities in complex reasoning tasks by emulating human-like reflective cognition.
-However, extending such capabilities to the multi-modal large language models (MLLMs) remains challenging due to the high cost of retraining vision-language alignments when upgrading the underlying reasoner LLMs.
-A straightforward solution is to decouple perception from reasoning, i.e., converting visual inputs into language representations (e.g., captions) that are then passed to a powerful text-only reasoner.
-However, this decoupling introduces a critical challenge: the visual extractor must generate descriptions that are both faithful to the image and informative enough to support accurate downstream reasoning.
-To address this, we propose Reasoning-Aligned Perceptual Decoupling via Caption Reward Optimization (RACRO) —a reasoning-guided reinforcement learning strategy that aligns the extractor’s captioning behavior with the reasoning objective.
-By closing the perception-reasoning loop via reward-based optimization, RACRO significantly enhances visual grounding and extracts reasoning-optimized representations.
-Experiments on multi-modal math and science benchmarks show that the proposed RACRO method achieves state-of-the-art average performance while enabling superior scalability and plug-and-play adaptation to more advanced reasoning LLMs without the necessity for costly multi-modal re-alignment.
+Recent breakthroughs in reasoning language models have significantly advanced text-based reasoning.
+On the other hand, Multi-modal Large Language Models (MLLMs) still lag behind, hindered by their outdated internal LLMs.
+Upgrading these is often prohibitively expensive, as it requires complete vision-language alignment retraining which is costly.
+To address this issue, we introduce Perception-Reasoning Decoupling, which modularizes the MLLM’s reasoning component and makes it easily replaceable.
+This approach redefines the MLLM’s role to convert multi-modal inputs into detailed textual outputs that can be processed by any powerful, external, text-only LLM reasoners.
+To align the MLLM’s perceptual output with the final reasoning task, we propose a novel reinforcement learning algorithm called Visual Perception Optimization (VPO).
+VPO rewards the MLLM based on the correctness of answers generated by the external reasoner to produce faithful and query-relevant captions.
+Together, this decoupling pipeline and VPO form our Reasoning-Aligned PerceptIon Decoupling (RAPID) approach.
+Empirical results show that RAPID achieves significant performance gains on multi-modal reasoning benchmarks.
+Crucially, RAPID enables a novel inference-time scaling paradigm: Once trained with VPO, the MLLM can be paired with any state-of-the-art LLM reasoner for consistent performance improvement without retraining.
 1 Introduction
-Recent developments in slow-thinking language models, such as OpenAI-o1 [26], DeepSeek-R1 [22], Gemini-Thinking [11], and Qwen3 [45], have led to significant advancements in the performance of large language models (LLMs) for complex math and science tasks.
-These models replicate human-like reflective thinking, intentionally modeling reasoning as a step-by-step process rather than relying on superficial shortcuts.
-As a result, they surpass fast-thinking models like GPT-4o [25] and Claude 3.5 [2], showing improvements of over 30% on mathematical benchmarks such as AIME24 [1] and AMC23 [15], along with approximately 10% gains on science benchmarks like GPQA [41].
-Motivated by these successes, there is growing interest in equipping multi-modal large language models (MLLMs) with similar slow-thinking capabilities to solve vision-based reasoning tasks that require both spatial perception and logical inference.
-A common strategy—adopted by methods like VL-Rethinker [47] and MM-EUREKA [38]—is to fine-tune a pre-trained MLLM (e.g., Qwen2.5-VL [5]) on multi-modal reasoning datasets using reinforcement learning with verifiable reward signals.
-However, this approach faces two critical limitations: (1) it is inherently restricted by the capabilities of underlying LLMs used for initialization; (2) adapting to more advanced LLMs necessitates redoing the vision-language alignment process, which involves burdensome multi-modal training that requires computing trillions of tokens.
-This raises a challenging research question: Can we efficiently integrate slow-thinking strategies into multi-modal reasoning without retraining the entire multi-modal LLM?
-A seemingly intuitive solution is to utilize a perceptual decoupling approach: separate the vision and reasoning components by converting multi-modal inputs into natural language captions, which are then passed to an external, more powerful text-only reasoner.
-While this idea promises flexibility and modularity, it introduces a fundamental challenge—the success of the reasoning process now critically depends on the quality of intermediate captions.
-If the visual extractor generates captions that are incomplete, overly generic, or hallucinatory, even the strongest reasoning LLMs will fail to give correct reasoning processes.
-As shown in Figure 3, existing MLLMs might struggle to produce reasoning-relevant descriptions, especially for complex mathematical and scientific visual inputs [51].
-Figure 1: Scalable multi-modal reasoning of RACRO. RACRO reveals better scalability compared to the traditional visual alignment, while remaining computationally efficient.
-To overcome this limitation, we propose Reasoning-Aligned Perceptual Decoupling via Caption Reward Optimization (RACRO) —a unified framework that not only decouples visual perception from reasoning, but also ensures their alignment through a reasoning-aware reinforcement learning procedure.
-In the proposed RACRO method, an MLLM-based extractor is responsible for encoding task-relevant visual information to natural language, while an external text-only LLM reasoner handles the slow-thinking reasoning process.
-Crucially, we introduce Caption Reward Optimization (CRO), a reinforcement learning algorithm that fine-tunes the extractor using reasoning outcomes as feedback.
-Given an input image and a question, the extractor generates multiple caption candidates; each is passed to the reasoner, and rewards are assigned based on whether the reasoner generates correct answers or not.
-This feedback loop aligns the extractor with the reasoning objective, guiding it to generate faithful and problem-relevant descriptions.
-By integrating perceptual decoupling with CRO into a unified framework, we enable effective and scalable slow-thinking in multi-modal reasoning systems.
-Experiments on challenging benchmarks—such as MathVerse [56], MathVision [48] and LogicVista [52]—show that the proposed RACRO method achieves state-of-the-art results.
-Furthermore, the proposed RACRO method supports plug-and-play adaptation to novel reasoning LLMs without extra multi-modal retraining, making it a practical and scalable solution for the rapid evolution of foundation models.
-It is illustrated in Figure 1, comparing the average performance of various MLLMs and their RACRO-enhanced variants across model sizes.
-For each RACRO-enhanced model group (e.g., Qwen2.5-VL-3B ($\bf{RACRO}$)), the original variant (leftmost, solid) is only trained with a base LLM reasoner with a small amount of data (e.g., 39K) using CRO, which can then achieve scalable performance with stronger LLMs (see detailed configurations on the choices of LLMs in Appendix A).
-Connecting these points yields a novel Pareto front that captures a superior trade-off between model size and performance than other base MLLMs.
-Our contributions can be summarized as the following three perspectives:
-  • We propose RACRO, a unified framework that decouples visual perception from reasoning and aligns them through reinforcement learning with verifiable rewards (RLVR).
-  • We introduce Caption Reward Optimization, a novel reinforcement learning strategy that uses reasoning success as a reward signal to optimize the extractor’s captioning behavior.
-  • We obtain state-of-the-art performance on multi-modal reasoning benchmarks, while maintaining scalability and eliminating the necessity for costly vision-language re-alignment.
-(a) Existing MLLM alignment methods
-(b) RACRO (ours)
-Figure 2: Comparisons between existing alignment methods for reasoning MLLMs and RACRO.
-For novel LLMs, existing methods (a) repeatedly conduct the intensive alignment process, while (b) RACRO decouples visual perception from reasoning (Sec.
-3.1) by learning to extract reasoning-aligned visual contexts with the proposed CRO paradigm (Sec.
+Recent reasoning language models, such as OpenAI-o1 (Jaech et al., 2024) and Qwen3 (Yang et al., 2025a), have driven significant gains in complex math and science tasks.
+By emulating a deliberate, step-by-step reasoning process akin to human reflection, these models avoid superficial shortcuts.
+As a result, they substantially outperform previous models like GPT-4o (Hurst et al., 2024), with improvements exceeding 30% on math benchmarks like AIME24 (AIME, 2024) and around 10% on science benchmarks like GPQA (Rein et al., 2024).
+Translating breakthroughs from the uni-modal text to the multi-modal domain remains a significant challenge.
+Existing multi-modal large language models (MLLMs), like Qwen2.5-VL (Bai et al., 2025), Gemma3 (Team et al., 2025a), and InternVL3 (Zhu et al., 2025), still struggle with reasoning and math-intensive tasks because their underlying LLMs are outdated or lack slow-thinking capabilities.
+While approaches like VL-Rethinker (Wang et al., 2025a) and MM-EUREKA (Meng et al., 2025) try to improve performance with reinforcement learning, their success is fundamentally restricted by the reasoning capability of the base LLM.
+The ideal solution, namely, switching the LLM with the most state-of-the-art one, is often prohibitive, as it requires repeating the entire, costly vision-language alignment process.
+This raises the critical question: Can we replace the LLM within an MLLM to unlock advanced reasoning [^1] In this paper, we focus on multi-modal math and science reasoning tasks.
+efficiently, without undertaking redundant vision-language retraining?
+To address that, we propose the Perception-Reasoning Decoupling pipeline, where we re-focus the MLLM’s primary role on perception.
+It first translates the multi-modal inputs into a comprehensive textual representation, which is then processed by a separate, powerful, external LLM for reasoning.
+This decoupling allows flexible alteration of the LLM reasoner, offering a path to circumvent the costly retraining cycle.
+Our key distinction from similar two-stage pipelines (Tiong et al., 2022; Guo et al., 2022; Hu et al., 2022; Gou et al., 2024; Lu et al., 2025) lies in the textual representation, which includes both a query-relevant caption and a tentative solution to ensure all essential visual information is captured for subsequent reasoning.
+However, the critical challenge in this new pipeline is that the generated textual outputs are not optimized for correct reasoning.
+Figure 1: Comparisons on multi-modal reasoning benchmarks on average performance and total model size between RAPID-enhanced Qwen2.5-VL series of models and the other existing MLLMs.
+Check the detailed numerical results in Appendix A and experimental settings in Sec.
+4.1.
+To overcome that, we introduce Visual Perception Optimization (VPO), a novel policy gradient algorithm operating through a reinforcement learning feedback loop where, given a user query, the MLLM first generates a group of query-relevant captions, which are then considered as contexts for the external LLM reasoner to generate final answers.
+With a correctness-based reward function, the perceptual MLLM is aligned with the reasoning objective, and guided to generate faithful and query-relevant captions optimized for the correctness of downstream reasoning.
+The combination of the Perception-Reasoning Decoupling pipeline with the VPO algorithm forms our overall approach, named Reasoning-Aligned PerceptIon Decoupling (RAPID)
+Empirically, RAPID achieves notable performance gains on challenging benchmarks such as MathVerse (Zhang et al., 2024), MathVision (Wang et al., 2024b) and LogicVista (Xiao et al., 2024).
+Moreover, as perception and reasoning are decoupled, the MLLMs trained with VPO generate textual outputs that can be directly fed to any LLM for reasoning.
+This eliminates the necessity for retraining, and enables RAPID to be a practical solution for the rapid evolution of MLLMs and reasoning LLMs.
+Figure 1 compares various MLLMs against the RAPID-enhanced Qwen2.5-VL series.
+For each RAPID-enhanced group (e.g., Qwen2.5-VL-3B), we optimize the MLLM with VPO using minimal data (39K).
+The resulting performance curves are generated simply by pairing the optimized MLLM with increasingly powerful external LLMs (see Appendix A for the choice of configurations), demonstrating a novel inference-time scaling paradigm.
+Our contributions can be summarized as follows:
+  • We introduce the Perception-Reasoning Decoupling pipeline, which redefines MLLMs’ focus to multi-modal perception, allowing the reasoning component to be flexibly replaced by any advanced external LLM without burdensome retraining.
+  • We propose Visual Perception Optimization (VPO), a novel policy gradient algorithm that aligns the MLLM’s perceptual outputs by using the correctness of the external LLM’s final answers with the perceptual outputs as contexts for reward signals.
+  • Combining both, RAPID achieves significant performance gains and introduces an efficient, novel “plug-and-play” inference-time scaling approach.
+By eliminating the costly retraining required by traditional methods, an one-time optimized MLLM can be paired with any stronger LLM for continual performance improvements, as demonstrated in Figure 1.
+(a) Existing MLLM alignment methods.
+(b) RAPID (ours).
+Figure 2: Comparisons between RAPID and existing alignment methods for reasoning MLLMs.
+For novel LLMs, existing methods (a) repeatedly conduct the compute-intensive alignment procedure, while (b) RAPID decouples the visual perception from text-only reasoning (Sec.
+3.1) by learning to extract reasoning-aligned visual contexts with the proposed VPO algorithm (Sec.
 3.2).
+Note that the caption penalty, as in Eq.
+3, is omitted here for simplicity.
+The flame and snowflake icons indicate the models are trainable and frozen, respectively, during the process.
 2 Related Work
-Multi-modal Large Language Models (MLLM).
-With the development of LLMs, integrating visual perception capability into LLMs has received increasing attention [19, 27, 30].
-To address the gap between different modalities, existing paradigms [5, 20, 21, 59] start with a text-only LLM and conduct pre-training on large-scale image captioning data to obtain a vision-language connector.
-Then, supervised fine-tuning (SFT) is applied with visual instruction tuning datasets.
-It is important to note that this procedure is computationally intensive.
-For example, Qwen2.5-VL [5] consumes around 4.1T tokens, and Kimi-VL [44], a more recent MLLM with MoE architecture, requires 2.3T tokens for the alignment.
-Under this paradigm, it remains formidable to upgrade MLLMs to newer LLMs.
-Instead, this paper proposes a computationally efficient solution for multi-modal reasoning.
-Empowering MLLMs with Slow-Thinking Capabilities.
-To empower MLLM with similar slow-thinking abilities, a widely adopted approach is to start with an existing MLLM (e.g., Qwen2.5-VL [5]) and perform reinforcement learning or knowledge distillation.
-For example, VL-Rethinker [47], MM-EUREKA [38] and NoisyRollout [31] apply GRPO [42] (or its variants [13, 31, 32, 38]) to MLLMs to learn deliberate reasoning patterns.
-Distillation-based methods, such as the OpenVLThinker [13], R1-OneVision [54], and Vision-R1 [24], perform supervised fine-tuning (SFT) on the vision-language reasoning data generated via a caption-and-reasoning pipeline [19], where visual inputs are converted to textual descriptions first, and strong text-only LLMs are then utilized to generate reasoning traces.
-However, their performances are limited by the base LLM (e.g., Qwen2.5 [53]), which often lags behind state-of-the-art reasoning models (e.g., DeepSeek-R1-Distilled [22]) in tasks like math.
-While replacing it with a stronger LLM is an intuitive solution, re-aligning vision and language through full retraining on trillions of tokens is prohibitively costly.
-In this work, we introduce the RACRO, a reasoning-aligned perceptual decoupling framework that bridges perception and reasoning via reward optimization, allowing seamless integration of stronger LLMs without retraining the vision-language interface, and meanwhile, demonstrating superior scalability w.r.t model sizes (c.f., Figure 1).
+Improving the Reasoning Ability of MLLMs.
+Start with an existing MLLM (e.g., Qwen2.5-VL (Bai et al., 2025)), a widely adopted approach is to perform reinforcement learning or knowledge distillation.
+For example, VL-Rethinker (Wang et al., 2025a), MM-EUREKA (Meng et al., 2025) and NoisyRollout (Liu et al., 2025a) apply GRPO (Shao et al., 2024) (or its variant (Liu et al., 2025b)) to MLLMs to learn deliberate reasoning patterns.
+Distillation-based methods, such as R1-OneVision (Yang et al., 2025b), Vision-R1 (Huang et al., 2025) and ReVisual-R1 (Chen et al., 2025) perform supervised fine-tuning (SFT) on reasoning data.
+However, both methods are restricted by the base LLMs (e.g., Qwen2.5 (Yang et al., 2024)), which lags behind state-of-the-art reasoning models (e.g., Qwen3-8B (Yang et al., 2025a)).
+While adopting a stronger LLM is an intuitive solution, re-aligning vision and language through full retraining on trillions of tokens is prohibitively costly.
+Caption-then-Reason Pipelines.
+To leverage LLM reasoning without intensive retraining, prior work explores similar “caption-then-reason" pipelines that decouple perception from reasoning.
+These approaches use Vision-Language Models (VLMs) (Radford et al., 2021; Li et al., 2022; Yu et al., 2022) or MLLMs for the perception task, while a separate LLM handles reasoning.
+Efforts in this area primarily focus on improving caption generation, for instance by selecting query-relevant image patches for captioning (Tiong et al., 2022; Guo et al., 2022), prompting MLLMs for query-aware captions (Gou et al., 2024), or enhancing captioning datasets (Hu et al., 2022; Lu et al., 2025).
+RAPID differs from these works in two key aspects.
+First, it includes a tentative solution in its generated output to better capture critical visual information.
+Second, while existing methods do not optimize captions for the final outcome, RAPID rewards the captioning process based on the correctness of the final answer produced by the reasoning LLM.
 3 Methodology
-In this section, we propose Reasoning-Aligned Caption Reward Optimization (RACRO), a unified framework that enables multi-modal slow-thinking reasoning by aligning the visual perception module with the downstream reasoning objectives.
-RACRO builds on the intuition of perceptual decoupling, where visual inputs are first translated into natural language representations that can be processed by powerful text-only reasoners.
-However, rather than treating the decoupling as static and assuming fixed captions suffice, RACRO closes the perception-reasoning loop through reinforcement learning, optimizing the captioning behavior of the visual extractor for downstream reasoning success.
-3.1 Reasoning-Aligned Perceptual Decoupling
-The perceptual decoupling paradigm is based on the assumption that the process of solving visual reasoning tasks can be separated into two distinct stages: visual perception and text-only reasoning.
-Specifically, given an image and a related question, if all relevant visual details can be losslessly converted into natural language, the problem can be solved using a text-only LLM.
-This decoupling allows for seamless upgrades to more powerful LLMs without the need to retrain or realign vision-language models.
-However, extracting all necessary visual details remains a severe challenge, as visual information can be dense, complex, and ambiguous [19, 51].
-To address that, we propose the Reasoning-Aligned Perceptual Decoupling framework, focusing on providing reasoning-aware visual details to the LLM instead of attempting to capture the entire visual input.
-Figure 2 gives an overview of the framework.
-Specifically, It consists of two steps:
-Figure 3: Inaccurate visual captions result in reasoning failure.
-MLLMs might produce hallucinated (left), vague, or incomplete captions (middle), particularly for complex visual reasoning, while captions containing necessary query-relevant contexts are essential for successful reasoning (right).
-Extraction.
-Guided by the goal of reasoning-aligned perceptual decoupling, we design an extraction process that generates visual details tailored to the reasoning needs for a given question.
-Instead of producing generic descriptions, we leverage an MLLM, such as Qwen2.5-VL [5] or InternVL3 [59], denoted as $M_{V}$, to produce two complementary outputs from an image $I$ and a query $\mathbf{q}$: a query-conditioned caption $\mathbf{c}=M_{V}(I,P_{\text{cap}}(\mathbf{q}))$, and a tentative solution $\hat{\mathbf{y}}=M_{V}(I,P_{\text{cot}}(\mathbf{q}))$, with each generated using distinct prompts (see Appendix B).
-The caption $\mathbf{c}$ captures visual content which is selectively relevant to the question, functioning as a reasoning-aligned abstraction.
-Although $\hat{\mathbf{y}}$ may not be fully accurate due to the limited reasoning capabilities of MLLMs, it often contains useful inductive signals that can inform and guide further downstream reasoning inference.
-Reasoning.
-The reasoning component is separately handled by a text-only large language model (LLM) (e.g., R1-Distilled-7B/32B [22], Qwen3-8B [45], or QwQ-32B [46]), denoted as $M_{R}$.
-It takes as input the question $\mathbf{q}$, the reasoning-aligned caption $\mathbf{c}$, and the tentative solution $\hat{\mathbf{y}}$, and produces the final answer: $\mathbf{y}=M_{R}(P_{\text{reason}}(\mathbf{q},\mathbf{c},\hat{\mathbf{y}}))$, where $P_{\text{reason}}$ is a prompt formatting function that integrates the inputs into a coherent textual reasoning context (see Appendix B).
-A core advantage of this decoupled framework is its flexibility: by separating reasoning from visual perception, we can boost performance on multi-modal tasks simply by upgrading to more capable LLMs for $M_{R}$, without modifying or retraining the MLLMs.
-While concurrent works have adopted similar decoupling paradigms—either as data engines for generating reasoning traces [24, 54] or as the inference framework [37]—they fail to investigate how to generate 1) query-conditioned and 2) reasoning-aligned visual representations, which is the focus of this paper in the following.
-3.2 Caption Reward Optimization
-While this design enables flexible upgrade of the reasoner without vision-language re-alignment, it exposes the system to a critical bottleneck: the quality of the caption directly affects the reasoning outcome.
-In practice, existing MLLMs may generate hallucinated [51], vague, or incomplete captions, particularly for complex visual reasoning tasks such as the geometric math questions (see Figure 3).
-To address this bottleneck, RACRO introduces Caption Reward Optimization—a reasoning-guided reinforcement learning algorithm that fine-tunes the captioning extractor using the feedback from a text-only reasoner, aiming at reinforcing captions that maximize likelihood of successful reasoning.
-Caption Reward Optimization
-is inspired by GRPO (Group Relative Policy Optimization) [42], which is a policy optimization algorithm originally developed to enhance the reasoning capability of text-only LLMs.
-In our setting, the policy $\pi_{\theta}$ to optimize becomes the MLLM $M_{V}$ performing visual captioning.
-For a given input pair $(I,\mathbf{q})$ of image and text question from the training set $p_{\mathcal{D}}$, the old policy generates $G$ caption rollouts, i.e., $\mathbf{c}\sim$ $\pi_{\theta_{\mathrm{old}}}(I,P_{\text{cap}}(\mathbf{q}))$.
-Denoting $R_{i}$ as the reward for the $i$-th rollout, the normalized advantage is $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ denotes the standard deviation of rewards within the group and the baseline reward is $\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$.
-The objective incorporates a surrogate loss clipped within $[1-\epsilon,1+\epsilon](\epsilon>0)$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown here) to stabilize optimization:
-$L(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\mathbf{c}\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{cap}}(\mathbf{q}))}$
-$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl{(}\frac{\pi_{\theta}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl{(}\frac{\pi_{\theta}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{c}_{i}\mid I,P_{\text{cap}}(\mathbf{q}))},\,1-\epsilon,\,1+\epsilon\Bigr{)}\hat{A}_{i}\Biggr{)}\Biggr{]}\textrm{.}$ (1)
-However, different from the math reasoning problem traditionally studied by GRPO, there are no extractable answers in the generated captions.
-Therefore, the widely adopted rule-based reward is no longer applicable to our problem.
-To tackle this, we propose to combine the rule-based reward function $r(\cdot,\cdot)$ and a reasoning LLM $M_{R}$ for reward computation.
-Specifically, the reward $R_{i}$ of a caption rollout $\textbf{c}_{i}$ is determined by whether a reasoner LLM $M_{R}$ could arrive at the ground-truth answer using this caption $\textbf{c}_{i}$ as contexts, which can be expressed as:
-$R_{i}=r(y_{\text{gt}},y_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(y_{i})),\quad\text{where}\ y_{i}=M_{R}(P_{\text{reason}}^{\prime}(\textbf{q},\textbf{c}_{i})),$ (2)
-where $y_{\text{gt}}$ denotes the ground-truth answer of a reasoning question and $\mathbbm{1}(\cdot)$ is an indicator function that outputs 1 if the final parsed prediction matches the ground-truth and 0 otherwise.
-Note that the reasoning process during reward computation is different from that in Sec.
-3.1 as the tentative response is not provided here (check the ablations on this configuration in Table 3).
-In summary, the proposed Caption Reward Optimization offers the following benefits:
-  • More accurate and task-relevant captions: As defined in Eq.
-(2), the reward encourages captions that facilitates successful downstream reasoning.
-Consequently, the MLLM learns to generate captions with less visual hallucination and more task-relevant visual details.
-  • Alignment scalability to any reasoners: Once the MLLM undergoes the Caption Reward Optimization process, even LLMs that are not involved during training can be incorporated into the reasoning framework because the captions are natural languages that generalize between different LLMs.
-This eliminates repeated alignments as done in the traditional MLLM instruction tuning approaches.
-As demonstrated in Figure.
-1, our RACRO reveals superior scalability during inference time, even though not been specifically trained for.
-Figure 4: Caption Reward Optimization reinforces captions that induce correct reasoning results via reinforcement learning with verifiable rewards (RLVR).
-4 Experiments
-4.1 Main Results
-Evaluation.
-Following VisualPRM [49], we evaluate our model on a diverse set of multi-modal reasoning benchmarks, including the MathVista (testmini) [36], MathVision (test) [48], MathVerse (vision-only) [56], MMMU (val) [55], WeMath [39], DynaMath [60], and the LogicVista [52].
-We report the worst-case accuracy for DynaMath and the strict accuracy for WeMath, while for the other benchmarks, we report the overall accuracy.
-All evaluations are conducted using VLMEvalKit [14].
-Table 1: Comparison on multi-modal reasoning benchmarks.
+This section describes the two main components of RAPID: perception-reasoning decoupling (Sec. 3.1) and visual perception optimization (Sec. 3.2). Figure 2 gives an overview of the approach.
+3.1 Perception-Reasoning Decoupling
+Given an image $I$ and a relevant query $q$, our perception-reasoning decoupling pipeline involves two consecutive stages: (i) Perception, where an MLLM (e.g., Qwen2.5-VL (Bai et al., 2025)) acts as a perception module to generate a group of textual outputs $O_{p}$ (detailed below) with respect to the image $I$ and a perception prompt.
+(ii) Reasoning, where a powerful text-only LLM reaonser (e.g., R1-Distilled-7B (Guo et al., 2025) or Qwen3-8B (Yang et al., 2025a)) receives the original query $q$ and a consolidated set of perceptual outputs, $O_{p}$, which are structured by a reasoning prompt $P_{r}$ (shown in Fig.
+12): $y=\operatorname{LLM}(P_{r}(q,O_{p}))$.
+A key advantage of this decoupling pipeline is that the textural outputs form a universal interface between perception (MLLMs) and reasoning (LLMs).
+This allows the reasoning LLMs to be upgraded independently, boosting performance without the necessity to retrain the MLLMs or alignment.
+A detailed empirical analysis is provided in Sec.
+4.3.
+Figure 3: Comparison of the strategies for visual perception $O_{p}$.
+Strategies for Visual Perception $O_{p}$.
+We explore strategies to generate precise perceptual outputs for reasoning:
+  • none: An empty set of perceptual outputs $O_{p}$, which serves as the control reference group.
+  • cap (Lu et al., 2025): A holistic image caption $o_{\text{cap}}=\operatorname{MLLM}(I,P_{\text{cap}})$ with the template $P_{\text{cap}}$ in Fig. 13.
+  • qcap (Gou et al., 2024): A query-relevant caption $o_{\text{qcap}}=\operatorname{MLLM}(I,P_{\text{qcap}}(q))$ with $P_{\text{qcap}}$ in Fig. 14.
+  • sol: A tentative solution $o_{\text{sol}}=\operatorname{MLLM}(I,P_{\text{sol}}(q))$ with the template $P_{\text{sol}}$ in Fig.
+15, which is “tentative” as it acts as contexts for LLMs instead of final answers.
+To evaluate how different compositions of the perceptual output set $O_{p}$ affect performance, we perform an experiment on seven multi-modal reasoning benchmarks (details in Sec.
+4.1 for details).
+In particular, we set $O_{p}$ to: (i) none; (ii) $\left\{o_{\text{cap}}\right\}$; (iii) $\left\{o_{\text{qcap}}\right\}$; (iv) $\left\{o_{\text{sol}}\right\}$; (v) $\left\{o_{\text{cap}},o_{\text{sol}}\right\}$; and (vi) $O_{p}=\left\{o_{\text{qcap}},o_{\text{sol}}\right\}$.
+The reasoning prompt $P_{r}$ (shown in Fig.
+12) is designed to provide a unified structure for all the above cases.
+We use Qwen2.5-VL-7B for perception and adopt Qwen3-8B and R1-Distilled-7B (refered to as R1-7B) for reasoning.
+Figure 3 shows the average accuracies obtained.
+As can be seen,
+  • Holistic captions outperform their query-relevant counterparts.
+This can be attributed to the MLLM’s extensive training on holistic image captioning tasks (Bai et al., 2025), whereas query-relevant captioning remains less optimized.
+However, with proper optimization (Sec.
+3.2), query-relevant captioning can offer an advantage by extracting contextually relevant visual details.
+  • Combining tentative responses and holistic captions achieves best results, delivering significant improvement (+7% w/ Qwen3-8B) over the original MLLM.
+This success is due to the complementary roles played by the caption and tentative response in reasoning.
+The caption provides the LLM with essential contexts for problem-solving, while the tentative response serves as a reference.
+While Figure 3 shows that cap+sol performs best, we will adopt qcap+sol as the default in the sequel.
+The reason is empirical.
+Our findings in Section 4.2 reveal that qcap+sol outperforms cap+sol once VPO (to be introduced in Section 3.2) is applied.
+This indicates that the query-relevant approach, while less optimized initially, possesses greater potential.
+3.2 Visual Perception Optimization (VPO)
+Although the combination of caption and tentative solution (i.e., both cap+sol and qcap+sol) demonstrates superior results in Figure 3, they are not optimized for the correctness of the final reasoning outcome.
+In other words, the MLLM generates its perception outputs without any feedback on whether these outputs actually guide the reasoning LLM to the correct answer.
+To address this limitation, we introduce Visual Perception Optimization (VPO).
+As illustrated in Figure 4, VPO establishes a reinforcement learning feedback loop that fine-tunes the MLLM for better captioning, explicitly rewarding it based on the correctness of the final answer produced by the reasoning LLM.
+Objective Design.
+Without the loss of generality, we describe VPO using the query-relevant caption (qcap) setting.
+VPO is inspired by Group Relative Policy Optimization (GRPO) (Shao et al., 2024), a policy optimization algorithm originally developed for text-only LLMs.
+In our setting, the policy $\pi_{\theta}$ to optimize is the MLLM performing visual captioning.
+For a given input pair $(I,q)$ from the training set $p_{\mathcal{D}}$, the old policy generates $G$ caption rollouts $\{o\sim\pi_{\theta_{\mathrm{old}}}(I,P_{\text{qcap}}(q))\}$ [^2] Here, we denote $o_{i}$ as the query-relevant captions qcap, rather than holistic captions cap or tentative solutions sol, as in Sec.
+3.1.
+Let $R_{i}$ be the reward for the $i$th rollout.
+The normalized advantage is $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ is the standard deviation of rewards within the group $R=\{R_{i}\}$ and $\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$ is the baseline reward.
+Thus, the objective of VPO, following the formulation of GRPO, can be represented as:
+$L(\theta)={}$ $\mathbb{E}_{(I,q)\sim p_{\mathcal{D}},o\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{qcap}}(q}))$
+$\Biggl[\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl(\frac{\pi_{\theta}(o_{i}\mid I,P_{\text{qcap}}(q))}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid I,P_{\text{qcap}}(q))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl(\frac{\pi_{\theta}(o_{i}\mid I,P_{\text{qcap}}(q))}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid I,P_{\text{qcap}}(q))},\,1-\epsilon_{\text{l}},\,1+\epsilon_{\text{h}}\Bigr)\hat{A}_{i}\Biggr)\Biggr]\textrm{,}$ (1)
+which incorporates a surrogate loss clipped to $1-\epsilon_{\text{l}},1+\epsilon_{\text{h}}](\epsilon_{\text{l}}>0,\epsilon_{\text{h}}>0))$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown) to stabilize optimization.
+Figure 4: Visual Perception Optimization (VPO) reinforces captions that induce correct reasoning results via reinforcement learning with verifiable rewards.
+Here we omit caption penalty for simplicity.
+Reward Design.
+GRPO is often used in math reasoning problems (Shao et al., 2024; Guo et al., 2025), in which the reward is determined by a simple rule because the model’s output is the final solution itself.
+However, in our setting, the MLLM generates an intermediate caption, from which the reasoning solution cannot be directly extracted.
+To address this, for each caption rollout $o_{i}$, we prompt the reasoning LLM to generate a final answer, and the reward is determined by whether this answer matches the ground-truth.
+This is formalized as follows:
+$\hat{R}_{i}=r(y_{\text{gt}},y_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(y_{i})),\quad\text{where}\ y_{i}=\operatorname{LLM}(P_{\text{reason}}^{\prime}(q,o_{i})),$ (2)
+where $y_{i}$ is the answer produced by the LLM from caption $o_{i}$, $\mathbbm{1}(\cdot)$ is the indicator function, and $P_{\text{reason}}^{\prime}(q,o_{i})$ (shown in Figure 16) is the reasoning prompt different from that used for inference (template in Figure 12) as it omits the tentative solution.
+The reward function $r(\cdot,\cdot)$ compares the predicted answer with the ground-truth $y_{\text{gt}}$.
+We ablate this design choice in Sec.
+4.2.
+During training, we observe reward hacking (details in Sec.
+4.2), where the MLLM directly solves the problem instead of performing captioning.
+Consequently, the model’s captioning ability does not improve.
+To address that, we impose a penalty on reward $\hat{R}_{i}$ if (i) $o_{i}$ leads to a correct answer; and (ii) $o_{i}$ does not contain a caption (determined by the policy MLLM $\pi_{\theta}$ via few-shot prompting):
+$R_{i}=\hat{R}_{i}-\lambda\,\mathbbm{1}\big(\hat{R}_{i}=1\land\neg\text{hasCap}(o_{i})\big),$ (3)
+where, $\lambda$ is a penalizing factor, and $\text{hasCap}(\cdot)$ checks if $o_{i}$ contains a caption (template in Figure 17).
+In summary, VPO offers two primary advantages:
+  • Generation of Reasoning-Aligned Captions: VPO uses the final reasoning outcome as a reward signal to optimize MLLMs, ensuring the captions are not merely descriptive but also functionally aligned for further reasoning.
+Check the quality improvement we demonstrate in Sec.
+4.4.
+  • LLM-Agnostic and Generalizable Improvement: VPO is an LLM-agnostic optimization, i.e., the optimized MLLMs communicate with the LLM reasoners via natural languages, and thus, the performance gains are generalizable across any LLMs.
+This enables a one-time alignment, which can be paired with any LLMs in a plug-and-play way without repeating VPO, as shown in Sec.
+4.3.
+In addition to captions, we further improve the quality of the tentative solution generated by the MLLM.
+As the tentative solution can be easily verified by a rule-based reward function, we apply GRPO on the MLLM.
+Details can be found in Appendix C.
+In the experiments, we optimize the MLLM with GRPO and VPO in a sequential manner, with VPO followed by GRPO.
+Table 1: Comparison on multi-modal reasoning benchmarks with respect to average accuracies. The best results are bold, while the second best are underlined. : short for GPT-OSS-120B-A5B.
 Model MathVista MathVision MathVerse MMMU WeMath DynaMath LogicVista AVG
@@
 GPT-4o-20241120 60.0 31.2 40.6 70.7 45.8 34.5 52.8 47.9
-o1 73.9 42.2 - 78.2 - - - -
 Open-Source Models
-InternVL3-2B-MPO 57.0 21.7 25.3 48.6 22.4 14.6 36.9 32.4
-OpenVLThinker-7B 70.2 25.3 - - - - - -
-R1-Onevision-7B 64.1 29.9 40.0 - - - - -
-Adora-7B 73.5 23.0 - - - - - -
-MM-Eureka-7B 73.0 26.9 - - - - - -
-InternVL3-8B-MPO 71.6 29.3 39.8 62.7 37.1 25.5 44.1 44.3
-InternVL3-14B-MPO 75.1 37.2 44.4 67.1 43.0 31.3 51.2 49.9
-MM-Eureka-32B 74.8 34.4 - - - - - -
-InternVL3-38B-MPO 75.1 34.2 48.2 70.1 48.6 35.3 58.4 52.8
-LLaVA-OV-72B 67.5 30.1 39.1 56.8 - - - -
-EMOVA-72B 69.9 - 50.0 59.7 - - - -
-VL-Rethinker-72B 80.3 43.9 - 68.8 - - - -
+$M$-Eureka-7B 73.0 27.9 46.1 54.9 34.7 22.6 48.3 43.9
+InternVL3-8B 73.6 29.3 39.8 62.7 37.1 25.5 44.1 44.6
+VL-Rethinker-7B 74.9 30.0 47.5 56.9 37.3 21.4 43.6 44.5
+ReVisual-R1-7B 70.8 43.0 52.7 55.7 40.7 30.5 51.2 49.2
+Ovis2-8B 71.8 25.9 42.3 57.4 27.2 20.4 39.4 40.6
+InternVL3-14B 75.1 37.2 44.4 67.1 43.0 31.3 51.2 49.9
+Ovis2-16B 73.7 30.1 45.8 60.7 45.0 26.3 47.4 47.0
+Gemma-3-12B 56.1 30.3 21.1 55.2 33.6 20.8 41.2 36.9
+$M$-Eureka-32B 74.7 36.6 51.5 62.0 37.1 33.5 58.2 50.5
+InternVL3-38B 75.1 34.2 48.2 70.1 48.6 35.3 58.4 52.8
+Ovis2-34B 76.1 31.9 50.1 66.7 51.9 27.5 49.9 50.6
+Gemma-3-27B 59.8 39.8 34.0 64.9 37.9 28.5 47.3 44.6
+QVQ-72B-Preview 70.3 34.9 48.2 70.3 39.0 30.7 58.2 50.2
 Qwen2.5-VL-72B 74.2 39.3 47.3 68.2 49.1 35.9 55.7 52.8
-QVQ-72B-Preview 70.3 34.9 48.2 70.3 39.0 30.7 58.2 50.2
-InternVL3-78B-MPO 79.0 43.1 51.0 72.2 46.0 35.1 55.9 54.6
+VL-Rethinker-72B 78.2 42.2 54.6 67.2 49.2 34.9 56.6 54.7
+InternVL3-78B 79.0 43.1 51.0 72.2 46.0 35.1 55.9 54.6
+Prior Caption-then-Reason Methods
+ECSO 64.6 42.7 42.8 61.4 38.4 25.0 39.4 44.9
+OmniCaptioner 67.5 43.3 48.0 62.2 38.7 30.5 56.2 49.5
+Qwen2.5-VL series and our RAPID-enhanced counterparts
 Qwen2.5-VL-3B 64.5 21.9 28.8 50.1 24.2 13.4 39.6 34.6
-w/ VisualPRM-8B-v1.1 69.5 35.1 36.3 56.0 39.2 10.6 43.4 41.4
-w/ RACRO (R1-7B) 64.5 (+0.0) 36.3 (+14.4) 39.3 (+10.5) 51.7 (+1.6) 31.5 (+7.3) 21.4 (+8.0) 46.5 (+6.9) 41.6 (+7.0)
-w/ RACRO (Qwen3-8B) 68.6 (+4.1) 39.6 (+17.7) 43.7 (+14.9) 58.6 (+8.5) 32.3 (+8.1) 29.3 (+15.9) 51.9 (+12.3) 46.3 (+11.7)
-w/ RACRO (R1-32B) 70.2 (+5.7) 41.5 (+19.6) 43.8 (+15.0) 58.4 (+8.3) 32.5 (+8.3) 30.3 (+16.9) 55.7 (+16.1) 47.5 (+12.9)
+w/ RAPID (Qwen3-8B) 69.6 40.8 48.6 60.9 39.1 29.3 56.4 49.2
 Qwen2.5-VL-7B 70.3 25.8 41.0 57.3 34.4 19.4 46.1 42.0
-w/ VisualPRM-8B-v1.1 74.6 27.1 45.0 56.1 46.3 17.8 48.5 45.1
-w/ RACRO (R1-7B) 71.9 (+1.6) 38.9 (+13.1) 45.3 (+4.3) 58.7 (+1.4) 37.1 (+2.7) 27.0 (+7.6) 49.7 (+3.6) 46.9 (+4.9)
-w/ RACRO (Qwen3-8B) 70.0 (-0.3) 41.5 (+15.7) 44.8 (+3.8) 60.6 (+3.3) 38.6 (+4.2) 26.5 (+7.1) 50.1 (+4.0) 47.4 (+5.4)
-w/ RACRO (R1-32B) 74.4 (+4.1) 42.5 (+16.7) 46.6 (+5.6) 61.1 (+3.8) 31.0 (-3.4) 30.3 (+10.9) 55.0 (+8.9) 48.7 (+6.7)
+w/ RAPID (Qwen3-8B) 76.1 43.7 52.2 64.7 45.4 32.7 57.7 53.2
 Qwen2.5-VL-32B 76.8 37.8 50.1 69.0 43.1 33.3 55.0 52.2
-w/ VisualPRM-8B-v1.1 77.7 39.9 52.2 69.6 50.4 33.1 53.9 53.8
-w/ RACRO (R1-7B) 75.4 (-1.4) 44.5 (+6.7) 52.9 (+2.8) 68.8 (-0.2) 48.3 (+5.2) 35.1 (+1.8) 54.6 (-0.4) 54.2 (+2.0)
-w/ RACRO (Qwen3-8B) 75.5 (-1.3) 47.1 (+9.3) 54.9 (+4.8) 67.9 (-1.1) 50.0 (+6.9) 35.7 (+2.4) 58.2 (+3.2) 55.6 (+3.4)
-w/ RACRO (R1-32B) 75.9 (-0.9) 46.7 (+8.9) 55.6 (+5.5) 69.0 (+0.0) 48.9 (+5.8) 37.1 (+3.8) 58.6 (+3.6) 56.0 (+3.8)
+w/ RAPID (Qwen3-8B) 76.8 47.0 54.4 67.8 48.5 36.5 60.4 55.9
+w/ RAPID (GPT-A5B) 75.9 52.1 54.3 69.8 50.8 38.3 60.4 57.4
+Qwen2.5-VL-72B 74.2 39.3 47.3 68.2 49.1 35.9 55.7 52.8
+w/ RAPID (GPT-A5B) 75.1 53.4 56.2 72.4 52.1 37.9 59.1 58.0
+4 Experiments
+4.1 Main Results
 Baselines.
-We compare our method with the following baselines:
-  • Proprietary models: GPT-4o [25], Claude-3.7-Sonnet [3], and Gemini-2.0-Flash [12] are general-purpose multi-modal models without dedicated reasoning optimization.
-In contrast, OpenAI o1 [26] is explicitly designed to enhance reasoning capabilities.
-These models represent the current proprietary state-of-the-art.
-  • Open-source general-purpose models: LLaVA-OneVision-72B (LLaVA-OV-72B) [27], EMOVA-72B [6], and the Qwen2.5-VL series [5] are primarily pretrained or instruction-tuned on large-scale image-text datasets, without explicit reasoning alignment.
-  • Open-source reasoning models: The InternVL3-MPO variants [59], OpenVLThinker-7B [13], R1-Onevision-7B [54], Adora-7B, MM-Eureka (7B/32B) [38], VL-Rethinker (7B/72B) [47], and QVQ-72B-Preview [40].
-These models are explicitly fine-tuned for reasoning tasks using methods such as SFT, DPO, and rule-based reinforcement learning.
-  • Multi-modal test-time scaling methods: We also compare our proposed method with multi-modal test-time scaling methods such as Best-of-N search (BoN) [49], which has been shown to effectively improve the reasoning accuracy of MLLMs.
-In particular, we consider the latest version, VisualPRM-8B-v1.1 [49].
-Note that in Table 1, BoN is adopted independently with RACRO.
-Implementation.
-To validate the effectiveness of the proposed method, we perform RACRO on the following MLLMs: Qwen2.5-VL series (3B, 7B, and 32B).
-During training, we use R1-Distilled-7B (referred to interchangeably as R1-7B) as the reasoner to compute reward signals for all MLLMs.
-During evaluation, in addition to R1-Distilled-7B, we also employ its 32B counterpart and Qwen3-8B to assess the scalability of our method.
-For training data, we adopt ViRL39K [47], a curated dataset of 38,870 verifiable multi-modal question-answer pairs tailored for reinforcement learning.
-We implement the caption reward optimization algorithm using the verl [43] framework with a global batch size of 256, a rollout temperature of 1.0, and a learning rate of $1\mathrm{e}^{-6}$.
-For each query, four captions are sampled.
-The clipping constant $\epsilon$ and KL-penalty coefficient $\beta$ are set to 0.2 and $1\mathrm{e}^{-3}$.
+We compare our method with the following baselines: (i) Proprietary models, including Claude-3.7-Sonnet (Anthropic, 2025), Gemini-2.0-Flash (DeepMind, 2025) and GPT-4o (Hurst et al., 2024); (ii) Open-source general-purpose MLLMs, including Qwen2.5-VL (3B/7B/32B/72B) (Bai et al., 2025), InternVL3 (8B/14B/38B/78B) (Zhu et al., 2025), Gemma-3 (12B/27B) (Team et al., 2025a) and Ovis2 (8B/16B/34B) (Lu et al., 2024b); (iii) Open-source MLLMs specialized for reasoning, including MM-Eureka (7B/32B) (Meng et al., 2025), VL-Rethinker (7B/72B) (Wang et al., 2025a), QVQ-72B-Preview (Qwen, 2024) and ReVisual-R1-7B (Chen et al., 2025).
+(iv) Latest Caption-then-Reason pipelines, such as ECSO (Gou et al., 2024) and OmniCaptioner (Lu et al., 2025).
+For fair comparisons, we use the GRPO-optimized MLLM (Qwen2.5-VL-7B) as captioner and Qwen3-8B as the reasoner.
+Evaluation
+is conducted on a diverse set of multi-modal reasoning benchmarks, e.g., MathVista (testmini) (Lu et al., 2024a), MathVision (test) (Wang et al., 2024b), MathVerse (vision-only) (Zhang et al., 2024), MMMU (val) (Yue et al., 2024), WeMath (Qiao et al., 2024), DynaMath (Zou et al., 2024), and LogicVista (Xiao et al., 2024).
+As in recent works  (Wang et al., 2025b; Zhu et al., 2025), we use VLMEvalKit (Duan et al., 2024) for evaluation, and report the worst-case accuracy for DynaMath, the strict accuracy for WeMath, and overall accuracy for the other benchmarks.
+Implementation Details of RAPID.
+We perform RAPID upon the Qwen2.5-VL series (3B, 7B, 32B, and 72B) MLLMs.
+During training, we use R1-Distilled-7B (R1-7B) as the reasoner to compute reward signals for all MLLMs.
+During evaluation, we adopt Qwen3-8B  [^3] We do not use R1-7B as we found the similar-sized Qwen3-8B performs better in Sec.
+4.3.
+(Yang et al., 2025a) and GPT-OSS-120B (Agarwal et al., 2025) as the LLM reasoners.
+For training data, we adopt ViRL39K (Wang et al., 2025a), a curated dataset of 38,870 verifiable multi-modal question-answer pairs tailored for multi-modal reasoning.
+We implement GRPO and VPO with verl (Sheng et al., 2025) with a global batch size of 256, a rollout temperature of 1.0, and a learning rate of $1\mathrm{e}^{-6}$.
+Implementation Details of GRPO.
+We set the number of rollouts to 8 for the 3B/7B MLLMs and 4 for the 32B/72B MLLMs.
+Following Yu et al.
+(2025), we remove KL regularization and use the "Clip-Higher" strategy, setting $\epsilon_{\text{l}}$ to $0.2$ and $\epsilon_{\text{h}}$ to $0.25$.
+When reporting performance with GPRO but without VPO (e.g., and in Table 2 or the baselines in Table 3), we select the best-performing checkpoints (with perception-reasoning decouple applied) at 400, 300, and 100 steps for the 3B, 7B, and 72B MLLMs, respectively, based on the average accuracies across the seven reasoning datasets (evaluated every 50 steps).
+GRPO is not applied to the 32B variant, as it has already been RL-tuned.
+Implementation Details of VPO.
+We set the number of rollouts to 4, KL-penalty coefficient $\beta$ to $\mathrm{e}^{-3}$, and penalizing constant $\lambda$ in Eq.
+(3) to $0.1$.
+VPO is applied to the MLLM following 200 steps of GRPO [^4] For the 3B model, we observe that training with VPO after GRPO results in slight forgetting of reasoning.
+To mitigate this, we switch back to GRPO optimization for an additional 100 steps after VPO.
+(except for the 32B model, which we directly use the original model).
+Similar to GRPO, we select the best checkpoints at 200, 150, 100, and 100 for the 3B, 7B 32B, and 72B models according to their average accuracies on the reasoning datasets.
 Results.
-Table 1 presents the performance of the proposed method.
-As shown, the average scores across seven datasets are significantly improved compared to the original MLLMs.
-For example, applying RACRO to Qwen2.5-VL-7B with a same-sized reasoner (R1-Distilled-7B) yields a +4.9% gain.
-Further improvements are observed when using a larger reasoner, such as a +6.7% increase with R1-Distilled-32B.
-Notably, when applying RACRO to Qwen2.5-VL-32B and evaluating with R1-Distilled-32B, we achieve state-of-the-art results on MathVision, MathVerse, and LogicVista, outperforming even larger models such as InternVL3-78B-MPO.
-Importantly, the Qwen3-8B and R1-32B reasoners are only used during inference time via a plug-and-play manner, since all trainings are conducted solely with the R1-7B reasoner.
-This demonstrates the scalability and generalization ability of our RACRO across both model sizes and (heterogeneous) reasoning LLM variants.
-Compared to Best-of-8 (Bo8) search, RACRO also achieves superior results.
-Notably, for stronger MLLMs such as Qwen2.5-VL-32B, VisualPRM-8B-v1.1 provides only limited improvement (e.g., +1.6%).
-RACRO, on the other hand, demonstrates greater scalability to stronger reasoners: when paired with a same-sized reasoner as VisualPRM-8B-v1.1, it improves the performance of Qwen2.5-VL-32B by +3.4%.
-This improvement is achieved without any re-training, highlighting RACRO’s plug-and-play compatibility with more capable reasoners.
-Table 2: Ablation study of different components of our RACRO.
-Sol. Cap. LLM CRO Math Vista Math Vision Math Verse MMMU We Math Dyna Math Logic Vista AVG
-31.5 36.0 10.0 39.2 11.4 3.8 22.8 22.1
-Qwen2.5-VL-3B
-64.5 21.9 28.8 50.1 24.2 13.4 39.6 34.6
-60.1 36.3 30.2 47.6 24.4 14.6 44.5 36.8
-59.5 36.2 12.7 47.8 25.0 13.4 35.1 32.8
-63.8 34.5 30.3 51.7 27.4 17.4 43.9 38.4
-64.5 36.3 39.3 51.7 31.5 21.4 46.5 41.6
-Qwen2.5-VL-7B
+Table 1 compares the performance of RAPID and baseline MLLMs on seven multi-modal reasoning datasets.
+It highlights two merits of RAPID: (i) It achieves significant performance gains on the reasoning tasks compared to the original MLLMs.
+For example, applying RAPID to Qwen2.5-VL-7B with a similar-sized LLM (Qwen3-8B) yields an average accuracy of $53.2\%$ (+11.2% compared to the original MLLM).
+Notably, when applying RAPID to Qwen2.5-VL-72B with GPT-OSS-120B as the LLM, we achieve the best average accuracy of $58\%$ across all the models compared, even surpassing proprietary MLLMs.
+(ii) RAPID achieves better performance-size trade-off.
+For example, Qwen2.5-VL-7B with RAPID (Qwen3-8B) with a total size of 15B outperforms larger models such as MM-Eureka-32B, InternVL3-38B and Ovis2-34B.
+Similarly, Qwen2.5-VL-32B with RAPID (Qwen3-8B) surpasses VL-Rethinker-72B and InternVL3-78B.
+Check Figure 1 for a better visualization of the performance-size trade-off.
+More analysis on the training and inference compute efficiency is provided in Appendix D.
+(iii) RAPID surpasses latest caption-then-reason methods (Gou et al., 2024; Lu et al., 2025), mainly due to the usage of tentative responses and VPO.
+Figure 5: General benchmark Results.
+Evaluation on General Benchmarks
+Although RAPID is specifically designed for multi-modal math and science reasoning, we verify that it does not hurt general abilities.
+We evaluated the VPO/GRPO-optimized Qwen2.5-VL-7B on general benchmarks in a “non-thinking” mode, per common protocols (Yang et al., 2025a; Wang et al., 2025c).
+As shown in Figure 5, its performance remains on par with the original model.
+This confirms that our method is a targeted enhancement for reasoning that preserves the model’s general abilities.
+(Benchmark details and Qwen2.5-VL-3B results are in Appendix E.1).
+Table 2: Ablation study of different components of RAPID (with Qwen2.5-VL-7B by default). VPO: VPO without the caption penalty; : using cap+sol for reasoning-perception decoupling.
+Decouple GRPO VPO Cap. penalty Math Vista Math Vision Math Verse MMMU We Math Dyna Math Logic Vista AVG
 70.3 25.8 41.0 57.3 34.4 19.4 46.1 42.0
-68.3 38.0 41.0 53.9 31.2 22.4 47.4 43.2
-54.4 35.8 27.9 50.6 23.7 7.6 34.9 33.6
-70.9 37.6 42.3 56.4 37.3 24.6 48.3 45.3
-71.9 38.9 45.3 58.7 37.1 27.0 49.7 46.9
-4.2 Ablation Study
-In this section, we first investigate the effectiveness of the proposed components, i.e., Reasoning-Aligned Perceptual Decoupling and Caption Reward Optimization.
-Then, we ablate the choices on reward computation during training.
+70.0 40.4 45.2 62.0 39.1 26.3 49.7 47.5
+72.7 43.2 50.0 63.3 41.1 28.7 54.1 50.5
+76.0 41.5 50.6 62.9 43.1 33.1 57.7 52.2
+76.1 43.7 52.2 64.7 45.4 32.7 57.7 53.2
+71.2 43.8 48.1 64.6 39.9 32.3 57.9 51.1
+74.2 42.5 50.8 62.0 39.4 31.9 56.6 51.1
+74.2 29.7 44.8 55.9 41.0 27.7 48.1 45.9
+75.0 29.8 42.0 55.8 40.8 23.0 46.3 44.7
+4.2 Ablation Studies
+In this section, we first investigate the effectiveness of the proposed components, i.e., reasoning-perception decoupling (denoted “Decouple") and VPO.
+For VPO, we ablate the choices on reward computation during training.
+Next, we assess the generalization and scalability of RAPID to different LLMs.
 We use the same training configurations as in Sec.
@@
 Qwen2.5-VL-3B/7B are adopted for ablations due to resource constraints.
-Unless otherwise specified, we use R1-Distilled-7B as the default LLM reasoner throughout the training and inference stages.
-Reasoning-Aligned Perceptual Decoupling & Caption Reward Optimization.
-For simplicity, we use the following shorthand markers to indicate the components:
-  • LLM indicates the usage of the text-only LLM reasoner $M_{R}$.
-  • Cap. denotes the use of the query-conditioned captions generated by the MLLM $M_{V}$.
-  • Sol. denotes the (tentative) solution provided by the MLLM $M_{V}$. Note that if the reasoner $M_{R}$ (i.e., LLM above) is not used, it is considered as the final output solution.
-  • CRO denotes the use of Caption Reward Optimization that optimizes the quality of the generated captions.
-We assess the RACRO framework by progressively enabling its components, as in Table 2.
-Starting from the baseline MLLM, we first introduce the tentative solution as the input to the LLM reasoner.
-This alone leads to notable performance improvements—e.g., from 34.6% to 36.8% for Qwen2.5-VL-3B and from 42.0% to 43.2% for the 7B model—demonstrating the reasoner’s ability to refine intermediate outputs from the MLLM.
-However, on tasks such as MathVista, MMMU, and WeMath, performance declines after introducing the LLM with tentative solutions only.
-To compensate for the lack of visual details, we incorporate query-conditioned captions alongside the tentative solution, which effectively restores the performance.
-Surprisingly, when captions are adopted alone without tentative solutions, benchmark performance drops significantly, suggesting that while captions can provide valuable visual context, they are often noisy—suffering from hallucinations and missing critical details.
-In contrast, the tentative solution acts as a reliable fallback for the reasoner when the visual signal is weak or ambiguous.
-To improve the quality of the query-conditioned captions, we apply the proposed Caption Reward Optimization strategy.
-As a result, we observe consistent performance gains across all tasks, confirming the quality issue of the captions and the effectiveness of the Caption Reward Optimization strategy.
-Table 3: Ablation studies on (a) reasoner types and (b) inputs for reward computation.
+Unless otherwise specified, we use R1-7B as the default LLM for training (reward computation) and Qwen3-8B for inference.
+Reasoning-Perception Decoupling & VPO.
+Table 2 presents a detailed ablation study of RAPID’s with the 7B MLLM (see Appendix E.2 for the 3B MLLM), which we analyze by incrementally adding each one to the baseline.
+Starting from the baseline MLLM (), we first apply perception-reasoning decoupling.
+This step alone () yields a significant 5.5% average improvement, demonstrating the immediate benefit of leveraging a stronger external LLM (Qwen3-8B) for reasoning.
+Building on this, we apply GRPO to enhance the MLLM’s perception by optimizing its tentative solutions, which adds another 3.0% to the average score ().
+We then apply VPO without the caption penalty (), and achieves a further 1.7% gain.
+Finally, incorporating the caption penalty leads to our full model (), adding another 1.0%.
+This brings the total improvement from our full VPO method to 2.7% over the model with only GRPO ().
+This caption penalty is crucial for VPO’s effectiveness, as it prevents reward-hacking where the model might generate solutions instead of the intended captions.
+Figure 9 confirms this: without the penalty, the ratio of rollouts containing valid captions diminishes rapidly, whereas with the penalty, it remains stable above 95% before 150 steps.
+Figure 6: Training dynamics of GRPO and VPO. Performance is evaluated under the perception-reasoning decouple pipeline.
+The results also highlight that GRPO and VPO are complementary, whose synergy is evident in two ways: 1) Removing either method from the full decoupled setup ( vs.
+, and vs.
+) results in suboptimal performance, confirming both are necessary.
+2) Figure 6 further visualizes this dynamic: after initial gains from GRPO begin to plateau, VPO provides a distinct secondary performance boost.
+Despite these gains, the decoupling strategy remains the most critical element.
+An MLLM improved by VPO and GRPO alone () still lags far behind the decoupled version (), underscoring its importance.
+We also note that VPO does not improve MLLM’s reasoning capabilities on its own ( vs.
+); the slight performance drop suggests minor forgetting during sequential training.
+Figure 7: Reward hacking without the caption penalty.
+Figure 8: Reward dynamics of adopting qcap and cap.
+Figure 9: Caption length trend with various LLMs.
+While holistic captions (cap+sol) initially outperform query-relevant ones (qcap+sol) as discussed in Section 3.1, this trend reverses after applying VPO (see and ).
+We hypothesize that this is because qcap is easier to optimize during VPO, as the query guides the MLLM to focus on relevant visual details.
+Instead, without such guidance, the MLLM struggles to identify pertinent information for cap.
+This is confirmed by Figure 9, which shows that training rewards for qcap increase steadily, while rewards for cap oscillate without consistent growth.
+Choices on Reward Computation.
+We study VPO reward computation by varying the reasoning LLM (during training) and its input content.
+We take the best-performing checkpoint (optimized with GRPO) as the baseline and evaluate under the perception-reasoning decoupled paradigm.
+First, we keep the input content as qcap+sol and test three LLMs with increasing reasoning capacity: Qwen2.5-7B-Instruct (weak), R1-7B (intermediate), and Qwen3-8B (strong), respectively.
+As shown in Table 3, the R1-7B LLM performs best.
+We hypothesize this is due to a trade-off in reasoning capacity, reflected in caption lengths.
+Figure 9 shows the caption lengths during training for various LLMs.
+We speculate that the stronger Qwen3-8B can succeed with more succinct captions,
+Table 3: Ablation studies on (a) LLM types and (b) input to LLM for reward computation.
 Configuration Q-VL-3B Q-VL-7B
-w/o CRO (baseline) 38.4 45.3
-(a) Reasoner Types
-Qwen-2.5-7B-Instruct 40.4 45.1
-R1-Distilled-1.5B 40.9 45.2
-R1-Distilled-7B 41.6 46.9
-(b) Reasoner Inputs
-Cap. & Sol. 40.8 45.8
-Cap. 41.6 46.9
-Choices on Reward Computation.
-Conducting CRO requires computing the reward of a caption rollout, which involves running the reasoner LLM.
-Here we discuss (i) what type of reasoner and (ii) what kind of inputs to the reasoner lead to more accurate reward computation.
-Specifically, for reasoners, we consider R1-Distilled-1.5B/7B and Qwen-2.5-7B-Instruct.
-For the reasoner input, we consider “Cap.”, which only includes the generated caption and “Cap.
-& Sol.” that further includes the MLLM’s tentative solution (pre-generated before training).
-Note that the later is exactly the configuration adopted in the proposed Perceptual Decouple framework.
-Table 3 shows the average performance with Qwen2.5-VL series as MLLMs (i.e., Q-VL-3B/7B) under different reasoner types and inputs.
-We observe that using smaller reasoners (i.e., R1-Distilled-1.5B) or models less specialized in reasoning (i.e., Qwen-2.5-Instruct) leads to worse overall performance.
-Intuitively, models performing poorly in reasoning are prune to generate incorrect solutions even though captions are accurate and comprehensive, providing noisy reward signals for CRO.
-However, larger reasoners significantly slow down the reward computing.
-Thus, we adopt R1-Distilled-7B as the reasoner as it strikes a balance between computational efficiency and reward accuracy by default.
-Regarding the input to the reasoner, we found passing only captions yields the best performance.
-Although “Cap.
-& Sol.” perfectly aligns with the Perceptual Decoupling framework during inference time, including a tentative response, allows the reasoner to take a shortcut that ignore the captions during reasoning, thereby producing a noisy reward signal for the captions.
-4.3 Analysis
-Scalability of CRO.
-In a practical scenario, after conducting CRO on an MLLM with a certain LLM (e.g., R1-Distilled-7B), it is desirable to switch to new LLMs (e.g., QwQ-32B) without additional training.
-We refer to this as the scalability of CRO and reveal that CRO indeed possesses this property.
-Figure 6 shows the overall task performance under the Perceptual Decoupling framework using different reasoners across varying model sizes and generations.
-The MLLMs adopted are from the Qwen2.5-VL series, both with and without CRO.
-As shown, although all MLLMs were optimized with CRO using R1-Distilled-7B, switching to different LLMs—whether in size (R1-Distilled-1.5B/32B) or generation (QwQ-32B)—still results in a notable performance boost over their non-CRO counterparts.
-This confirms the scalability of CRO.
-Furthermore, we observe that this performance gain becomes more pronounced with larger or stronger LLMs.
-This is because, without CRO, these models are significantly bottlenecked by inaccurate or low-quality visual captions, preventing them from leveraging their full reasoning potential.
-In contrast, smaller or weaker LLMs are limited not only by caption quality but also by their own reasoning capacity, making the relative improvement from CRO less substantial.
-This highlights the effectiveness of CRO in improving the quality of image captions.
-Further, with stronger LLMs, the performance gap between 3B and 7B gradually diminishes.
-This reveals that reasoning is better decoupled from perception.
-Figure 5: Ablation on the choice of reasoner.
-Figure 6: Ablation on CRO and GRPO.
-CRO and GRPO are Complementary.
-The proposed CRO objective effectively enhances the quality of query-conditioned captions, thereby improving overall performance.
-Another direction for improving the reasoning capability of the LLM is to produce more accurate tentative responses, which can be achieved through reinforcement learning methods such as GRPO.
-To evaluate their contributions, we apply CRO and GRPO separately to the original MLLMs (e.g., Qwen2.5-VL) and compare their average task performance under the perceptual decoupling framework (details on GRPO are in Appendix C.1).
-To investigate whether the two objectives are complementary, we also assess a combined setting—denoted as “CRO+GRPO”—where tentative responses and captions are generated by models fine-tuned separately with GRPO and CRO.
-For comparison, results from models without additional training are labeled as “No Training”.
-Additionally, we further include VL-Rethinker-7B [47], which is initialized from Qwen2.5-VL-7B and thoroughly fine-tuned using GRPO.
-We further fine-tune this model with CRO and report the resulting performance as “GRPO+CRO”.
-All results are visualized in Figure 6.
-As shown, both CRO and GRPO improve the performance under the perceptual decoupling framework.
-More importantly, for all MLLMs, combining both objectives brings notable improvement over the individuals.
-This suggests that the proposed CRO is complementary to GRPO.
-Note that we did not report results with GRPO training in Table 1 to highlight the impact of perceptual decoupling.
-Figure 7: Pairwise comparison on caption quality.
-Evaluating the Quality of Captions.
-To see whether the quality of the generated captions are improved, we conduct pairwise comparisons [9, 34] among captions generated by Qwen2.5-VL-3B trained with and without CRO.
-We randomly sample 1000 captions for each evaluation dataset and use GPT-4o  [^1] https://genai.ust.hk/ as the judge to determine the preferred caption.
-The judge receives the image, the question alongside two captions, and decides the winner or a tie game.
-We instruct the judge to choose captions that include more comprehensive and accurate details required to answer the question and exclude any solving process in the captions (the detailed prompt is in Appendix C.2).
-Furthermore, we alternate the order of the two captions to avoid position bias following [57].
-Results are reported in Figure 7.
-It can be observed that across all datasets, models trained with CRO offer a clear advantage over the original ones.
-This confirms the quality issues of captions discovered by this paper and highlights the effectiveness of CRO.
-To ensure robustness of our automated evaluation, we further conduct human assessment through the same pairwise comparison framework with details in Appendix C.2.
-The results reinforces the quality improvement of the captions.
+w/o VPO (baseline) 46.1 50.5
+(a) LLM Types
+Qwen2.5-7B-Instruct 47.5 51.9
+R1-Distilled-7B 47.9 53.2
+Qwen3-8B 47.1 51.3
+(b) LLM Input Contents
+qcap+sol 47.4 49.5
+qcap 47.9 53.2
+thus inadvertently rewarding brevity, while the weaker Qwen2.5-Instruct-7B incentivizes informative captions, causing length to surge. R1-7B strikes an effective balance, making it our default choice.
+Next, we examine the choice of input content: using caption alone (qcap) versus using the caption plus a tentative solution (qcap+sol).
+As in Table 3, using only the caption is superior.
+Including tentative solutions allows the LLM to take a shortcut during training—relying on the solutions while ignoring captions—which generates a noisy reward signal ineffective for optimizing caption quality.
+[^5] Notably, this optimal input format differs from that used in the perception-reasoning decoupling stage.
+Figure 10: Performance with different LLMs. (Only R1-7B is used in training)
+4.3 Generalization and Scaling with Different LLMs.
+A practical requirement is that our MLLM, once optimized, should generalize to new, unseen LLMs at inference time without retraining.
+To test this, we perform VPO on the GRPO-trained MLLM using only R1-7B as the LLM for this optimization step.
+We then evaluate its performance against the baseline (the MLLM without VPO) by pairing both MLLMs with a diverse range of different LLMs at inference time, as shown in Figure 10.
+.
+We have three observations.
+First, the performance gain from VPO generalizes effectively.
+The gap between the VPO-trained model (solid curves) and the baseline (dashed curves) is maintained or widened when using stronger LLMs, revealing that the benefit is not confined to the specific LLM used for training.
+Second, the RAPID’s scalability is evident as absolute performance trends upward when using more capable LLMs, although this improvement is not strictly monotonic with model size.
+Additionally, among the LLMs tested, Qwen3-8B strikes the best balance between performance and model size, establishing it to be our default choice for the inference stage.
+Note that the optimal LLMs for training and inference might differ (c.f., Table 3 and Figure 10).
+Figure 11: Pairwise comparison on the caption quality with and without VPO.
+4.4 Evaluation on Caption Quality
+We validate the improved quality of the captions generated by VPO-optimized MLLMs via a pairwise comparisons (Chen et al., 2023; Liu et al., 2024b) using GPT-4o (OpenAI, 2024) as a judge.
+With 1000 random samples per dataset, GPT-4o compares captions from Qwen2.5-VL-3B trained with and without VPO.
+The judge is instructed to prefer captions with more comprehensive and accurate details required to answer the question, while excluding any solving process (the prompt is in Appendix E.3).
+We alternate the caption order to mitigate position bias (Zheng et al., 2023).
+As shown in Figure 11, the VPO-optimized MLLM’s captions demonstrate a clear advantage across all datasets, highlighting the VPO’s effectiveness (We conduct a case study on caption quality in Appendix G).
+Moreover, we validate these findings with human assessment in Appendix E.3.
 5 Conclusion
-In this paper, we propose the RACRO framework to construct multi-modal reasoning models efficiently.
-By decoupling multi-modal reasoning with the visual perception and text-only reasoning, RACRO embraces the extraordinary reasoning capabilities of frontier LLMs, while getting rid of the burdensome visual re-alignment.
-Enhanced with Caption Reward Optimization, precise captions are reinforced to provide abundant visual contexts for better downstream reasoning, effectively scaling towards more advanced LLM reasoners during inference time.
-We obtain state-of-the-art performance on multiple multi-modal reasoning benchmarks, while remaining computationally efficient.
-Appendix
+This paper proposes RAPID, an efficient method for constructing multi-modal reasoning models.
+By decoupling visual perception (MLLM) from text-only reasoning (LLM), RAPID leverages the advanced reasoning of frontier LLMs while avoiding burdensome visual re-alignment.
+Enhanced with Visual Perception Optimization, this method reinforces precise captions to provide rich visual context, improving reasoning and effectively scaling to more advanced LLMs at inference time.
+Our approach achieves significant accuracy gains on multiple multi-modal reasoning benchmarks while remaining computationally efficient.
 Appendix A Model Configuration for Figure 1
-For each RACRO-enhanced model group (e.g., Qwen2.5-VL-3B (RACRO)) in Figure 1, we train the original model (e.g., Qwen2.5-VL-3B, they are shown as hollow squares with different colors in the figure) using both the proposed CRO and the existing GRPO objectives, where the former encourages the MLLM to generate query-conditioned captions with higher quality while the later optimizes it to give better reasoning traces.
-The details for CRO and GRPO can be found in Sec.
-4.1 and Appendix C.1.
-We then pair the trained MLLM with different LLM reasoners under our RACRO framework.
-Table 4 shows the configuration for each RACRO-enhanced model group.
-Specifically, for each model in the group, we report the total model size (B), paired LLMs, data used to conduct GRPO and CRO and the average performance across 7 tasks.
-Note that the results for Qwen2.5-VL-32B group does not involve GRPO training because we obtain worse results than the original model after GRPO.
-Results for other MLLMs without RACRO are directly taken from Table 1.
-Table 4: Model Configurations for each RACRO-enhanced model group in Figure 1. “-” denotes the corresponding item (e.g., LLM reasoner, GRPO/CRO training) is not applied.
-Size (B) LLM GRPO CRO Avg. Performance
+For each RAPID-enhanced model group (e.g., Qwen2.5-VL-3B (RAPID)) in Figure 1, we train the original model (e.g., Qwen2.5-VL-3B) using both the proposed VPO and GRPO objectives, where the former encourages the MLLM to generate query-relevant captions with higher quality while the later optimizes it to give better reasoning traces.
+The details for VPO and GRPO can be found in Sec.
+4.1 and Appendix C.
+We then pair the trained MLLM with different LLMs under our RAPID.
+Table 4 shows the configuration for each RAPID-enhanced model group.
+Specifically, for each model in the group, we report the total model size (B), paired LLMs, data used to conduct GRPO and VPO and the average performance across 7 tasks.
+Note that the results for Qwen2.5-VL-32B group does not involve GRPO training because it has already undergone an RL stage before release (Bai et al., 2025).
+Results for other MLLMs are consistent with Table 1.
+Table 4: Model Configurations for each RAPID-enhanced model group in Figure 1. “-” denotes the corresponding item (e.g., LLM reasoner, GRPO/VPO training) is not applied.
+Size (B) LLM GRPO VPO Avg. Performance
 Qwen2.5-VL-3B
-3 - - - 34.6
-7 Qwen3-4B ViRL39K ViRL39K 45.4
-11 Qwen3-8B ViRL39K ViRL39K 46.3
-35 R1-32B ViRL39K ViRL39K 47.8
+7 Qwen3-4B ViRL39K ViRL39K 47.3
+11 Qwen3-8B ViRL39K ViRL39K 49.2
+17 Qwen3-14B ViRL39K ViRL39K 50.1
 Qwen2.5-VL-7B
-7 - - - 44.5
-11 Qwen3-4B ViRL39K ViRL39K 48.9
-15 Qwen3-8B ViRL39K ViRL39K 50.2
-39 R1-32B ViRL39K ViRL39K 50.8
+11 Qwen3-4B ViRL39K ViRL39K 51.0
+15 Qwen3-8B ViRL39K ViRL39K 53.2
+21 Qwen3-14B ViRL39K ViRL39K 53.6
 Qwen2.5-VL-32B
-32 - - - 52.2
-39 R1-7B - ViRL39K 54.2
-40 Qwen3-8B - ViRL39K 55.6
-64 R1-32B - ViRL39K 56.0
-Appendix B Prompt Templates Used in RACRO
-In Table 5, we provide an index of the prompt templates used in RACRO.
-Table 5: Index of Prompt templates for RACRO.
+36 Qwen3-B - ViRL39K 54.6
+40 Qwen3-8B - ViRL39K 55.9
+152 GPT-OSS-120B - ViRL39K 57.4
+Appendix B Prompt Templates
+In Table 5, we provide an index of the prompt templates used in RAPID.
+Table 5: Index of Prompt templates for RAPID.
 Component Purpose Notation Prompt Template
-MLLM Tentative response $P_{\text{cot}}$ Figure 10
-MLLM Query-conditioned captions $P_{\text{cap}}$ Figure 11
+MLLM Holistic captions $P_{\text{cap}}$ Figure 13
+MLLM Query-relevant captions $P_{\text{qcap}}$ Figure 14
+MLLM Tentative response $P_{\text{sol}}$ Figure 15
+MLLM Caption Penalty - Figure 17
 LLM Reasoner Inference $P_{\text{reason}}$ Figure 12
@@
 Figure 12: Prompt templates used by the reasoner LLM for inference.
+Figure 13: Prompt templates used by the MLLM to obtain the holistic captions.
+Figure 14: Prompt templates used by the MLLM to obtain the query-relevant captions.
+Figure 15: Prompt templates used by the MLLM to obtain the tentative response. The placeholder is for the question.
 Figure 16: Prompt templates used by the reasoner LLM for training.
-Figure 10: Prompt templates used by the MLLM to obtain the tentative response. The placeholder is for the question.
-Figure 11: Prompt templates used by the MLLM to obtain the query-conditioned captions.
-Appendix C Experimental Details
-C.1 Details for GRPO Training
-GRPO
-(Group Relative Policy Optimization) [42] is a policy optimization algorithm originally developed to enhance the reasoning capability of text-only LLMs.
+Figure 17: Prompt templates used by the MLLM for caption panelty.
+Figure 18: Prompt templates used for GPT evaluations on caption qualities.
+Appendix C Formulations of GRPO
+GRPO (Shao et al., 2024) is a policy optimization algorithm originally developed to enhance the reasoning capability of text-only LLMs.
 In our setting, the policy $\pi_{\theta}$ to optimize becomes the MLLM.
-For a given input pair $(I,\mathbf{q})$ of image and text question from the training set $p_{\mathcal{D}}$, the old policy generates $G$ rollouts, i.e., $\mathbf{o}\sim$ $\pi_{\theta_{\mathrm{old}}}(I,P_{\text{cot}}(\mathbf{q}))$.
+For a given input pair $(I,q)$ of image and text question from the training set $p_{\mathcal{D}}$, the old policy generates $G$ rollouts, i.e., $o\sim$ $\pi_{\theta_{\mathrm{old}}}(I,P_{\text{sol}}(q))$.
 Denoting $R_{i}$ as the reward for the $i$-th rollout, the normalized advantage is $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ denotes the standard deviation of rewards within the group and the baseline reward is $\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}$.
-The objective incorporates a surrogate loss clipped within $[1-\epsilon,1+\epsilon](\epsilon>0)$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown here) to stabilize optimization:
-$L(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\mathbf{o}\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{cot}}(\mathbf{q}))}$
-$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))}{\pi_{\theta_{\mathrm{old}}}(\mathbf{o}_{i}\mid I,P_{\text{cot}}(\mathbf{q}))},\,1-\epsilon,\,1+\epsilon\Bigr{)}\hat{A}_{i}\Biggr{)}\Biggr{]}\textrm{.}$ (3)
-The reward $R_{i}$ for the $i$-th rollout is expressed as: $R_{i}=r(y_{\text{gt}},\textbf{o}_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(\textbf{o}_{i})),$ where $y_{\text{gt}}$ denotes the ground-truth answer of a reasoning question and $\mathbbm{1}(\cdot)$ is an indicator function that outputs 1 if the final parsed prediction matches the ground-truth and 0 otherwise.
-Implementation Details.
-Similar to CRO, for training data, we adopt ViRL39K [47] and implement GRPO using the verl [43] framework with a global batch size of 256, a rollout temperature of 1.0, and a learning rate of $1\mathrm{e}^{-6}$.
-For each query, eight rollouts are sampled.
-The clipping constant $\epsilon$ and KL-penalty coefficient $\beta$ are set to 0.2 and $1\mathrm{e}^{-3}$, respectively.
-C.2 Details for Pairwise Comparisons
+The objective incorporates a surrogate loss clipped within $1-\epsilon,1+\epsilon](\epsilon>0)$ and a KL-penalty $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ weighted by $\beta$ (not shown here) to stabilize optimization:
+$L(\theta)={}$ $\mathbb{E}_{(I,q)\sim p_{\mathcal{D}},o\sim\pi_{\theta_{\text{old}}}(\cdot\mid I,P_{\text{sol}}(q))}$
+$\Biggl[\frac{1}{G}\sum_{i=1}^{G}\min\ \!\Biggl(\frac{\pi_{\theta}(o_{i}\mid I,P_{\text{sol}}(q))}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid I,P_{\text{sol}}(q))}\hat{A}_{i},\mathrm{clip}\ \!\Bigl(\frac{\pi_{\theta}(o_{i}\mid I,P_{\text{sol}}(q))}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid I,P_{\text{sol}}(q))},\,1-\epsilon,\,1+\epsilon\Bigr)\hat{A}_{i}\Biggr)\Biggr]\textrm{.}$ (4)
+The reward $R_{i}$ for the $i$-th rollout is expressed as: $R_{i}=r(y_{\text{gt}},o_{i})=\mathbbm{1}(y_{\text{gt}}=\operatorname{parse}(o_{i})),$ where $y_{\text{gt}}$ denotes the ground-truth answer of a reasoning question and $\mathbbm{1}(\cdot)$ is an indicator function that outputs 1 if the final parsed prediction matches the ground-truth and 0 otherwise.
+Appendix D More Analysis
+D.1 Analysis on the Training Compute Efficiency of RAPID
+RAPID’s decoupled design enables the flexible adoption of recent LLM reasoners, such as Qwen3 (Yang et al., 2025a), without retraining.
+This raises a critical question: how does our efficient approach compare against models that require full and costly retraining of their visual-language alignment to integrate the latest LLMs?
+To investigate this trade-off between performance and computational cost, we compare RAPID with two leading MLLMs also built on the Qwen3-8B LLM: Keye-VL (Team et al., 2025b) and InternVL3.5 (Wang et al., 2025c).
+Table 6 presents a comparative analysis, reporting average accuracy across seven multi-modal reasoning tasks alongside the training tokens and estimated training FLOPs (calculated as model size × training tokens).
+The number of training tokens for Keye-VL-8B and InternVL3.5-8B are sourced from their respective technical reports.
+As can be seen, although still inferior to the end-to-end methods, RAPID with Qwen2.5-VL-7B can achieve $90.8\%$ of Keye-VL-8B performance with $1250\times$ less training FLOPs, and $88.2\%$ of InternVL3.5-8B performance with $864.2\times$ training cost reduction.
+Thanks to the remarkable training efficiency, we can adopt larger MLLMs such as Qwen2.5-VL-32B, we achieve comparable ($92.7\%$) performance with InternVL3.5-8B but with $1025\times$ less training FLOPs
+Table 6: Training cost comparison. As we did not apply GRPO to Qwen2.5-VL-32B, it consumes less training tokens (100M) than Qwen2.5-VL-7B (550M).
+Method AVG Accuracies # Tokens FLOPs (Ratio)
+Keye-VL-8B (Team et al., 2025b) 58.6 600B $1500\times$
+InternVL3.5-8B (Wang et al., 2025c) 60.3 410B $1025\times$
+Qwen2.5-VL-7B w/ RAPID (Qwen3-8B) 53.2 550M $1.2\times$
+Qwen2.5-VL-32B w/ RAPID (Qwen3-8B) 55.9 100M $1\times$
+D.2 Analysis on the Inference Compute Efficiency of RAPID
+To evaluate the inference compute efficiency of RAPID, we estimate the computational cost across the seven evaluation datasets from Table 1.
+For each dataset, we perform inference on a sample of 100 examples and approximate the compute as model size $\times$ number of generated tokens.
+For our staged RAPID approach, the total compute is the sum from the perception and reasoning stages, with the compute for each stage calculated using its respective model size.
+We benchmark these results against top-performing MLLMs from Table 1.
+Figure 19 plots the resulting average accuracy against inference compute for both RAPID-enhanced models and their baselines.
+The analysis reveals two key findings:
+  • RAPID achieves a favorable trade-off between accuracy and inference compute.
+This efficiency is highlighted by specific configurations that demonstrate Pareto-optimality.
+For instance, Qwen2.5-VL-7B w/ RAPID (R1-7B) provides a better accuracy-to-compute ratio than ReVisual-R1-7B.
+In another example, Qwen2.5-VL-7B w/ RAPID (Qwen3-8B) outperforms the much larger Qwen2.5-VL-72B while being more computationally efficient.
+  • RAPID demonstrates strong scalability with inference compute.
+The architecture is designed such that allocating more computational resources at inference—for example, by swapping in a more powerful LLM reasoner—consistently yields higher accuracy.
+Figure 19: Inference compute versus average accuracy.
+Appendix E More Experimental Details
+E.1 Evaluation on General Benchmarks
+Unlike the specialized math and science reasoning benchmarks, our general evaluation suite is chosen to assess foundational vision-language capabilities.
+We select MME (Fu et al., 2024a), MMBench-v1.1 (Liu et al., 2024a),
+Figure 20: General benchmark Results. (Qwen-2.5-VL-3B)
+MM-Vet (Yu et al., 2023), SEED-Image [^6] We evaluated on the “Image” split.
+(Li et al., 2023), MMstar (Chen et al., 2024), RealWorld-QA (RW-QA) (xAI, 2024), MMT-Bench (MMT) (Ying et al., 2024), and BLINK (Fu et al., 2024b), which cover tasks such as object recognition, text recognition (OCR), spatial awareness and so on.
+Figure 20 presents the results for Qwen2.5-VL-3B.
+Similar to the observations for the 7B model (Figure 5), the VPO/GRPO-optimized model performs comparably to the original MLLM (Note we do not report results with Qwen2.5-VL-32B/72B as they show the same observations).
+This confirms that RAPID preserves general-purpose abilities across different model scales.
+E.2 Ablation Study on Qwen2.5-VL-3B
+In this section, we extend our ablation study to the smaller Qwen2.5-VL-3B model, with results presented in Table 7.
+While the results are largely consistent with those from its 7B counterpart (Table 2), a critical difference emerges.
+The 3B model necessitates an additional GRPO stage () following VPO to restore its reasoning capabilities [^7] We hypothesize that VPO degrades the quality of the intermediate reasoning steps passed to the LLM, an effect not always visible in the final accuracy., which in turn yields considerable accuracy gains.
+We attribute this requirement to the limited capacity of the 3B model, where optimizing for the VPO task appears to degrade its inherent reasoning performance—a trade-off that is less pronounced in the larger 7B model.
+Table 7: Ablation study of different components of RAPID (with Qwen2.5-VL-3B).
+VPO: VPO without the caption penalty; : using cap+sol for reasoning-perception decoupling.
+: After VPO, we additionally conduct GRPO to recover its reasoning ability.
+Decouple GRPO VPO Cap. penalty Math Vista Math Vision Math Verse MMMU We Math Dyna Math Logic Vista AVG
+64.5 21.9 28.8 50.1 24.2 13.4 39.6 34.6
+65.5 39.1 31.9 59.0 31.1 24.8 48.3 42.8
+68.5 40.0 39.2 61.0 35.1 26.9 51.7 46.1
+68.5 39.4 43.4 59.9 37.4 27.9 55.3 47.4
+69.0 39.7 44.3 60.9 38.6 27.3 55.3 47.9
+69.6 40.8 48.6 60.9 39.1 29.3 56.4 49.2
+67.0 40.9 44.3 58.0 33.2 28.9 54.4 46.7
+68.8 41.0 43.8 59.8 34.8 28.7 54.4 47.3
+69.1 26.9 38.2 56.9 34.0 20.0 42.5 41.1
+68.8 26.9 39.8 49.4 33.0 21.6 46.3 40.8
+E.3 Details for Pairwise Comparisons
 Prompt for GPT evaluations.
@@
 Consistent with the details in Sec.
-4.3, this instructs the GPT to (1) choose captions that include more comprehensive and accurate details required to answer the question and (2) exclude any solving process in the captions.
-Figure 12: Prompt templates used for GPT evaluations on caption qualities.
+4.4, this instructs the GPT to (1) choose captions that include more comprehensive and accurate details required to answer the question and (2) exclude any solving process in the captions.
 Human evaluations.
 We present the details of the human evaluation conducted for the pairwise comparison experiment.
-For this, 100 questions are randomly sampled from the testmini set of MathVista, and captions are generated using Qwen2.5-VL-3B, trained with and without CRO.
-A total of 4 trained human annotators are recruited, with each annotator responsible for evaluating 25 questions to determine a winner or a tie.
-The annotators’ primary task is to evaluate a question and its two anonymous model-generated answers, then select the one they deem of higher quality, following the scoring principle adopted from [4].
-If neither answer stood out, they are instructed to choose "tie."
-In Table 6, we report the win/tie/lose ratio (i.e., “win” means captions generated by MLLMs with CRO is better) and an additional measure of consistency, calculated by the agreement rate between GPT-4o and human judgments.
-As can be seen, Qwen2.5-VL-3B trained with CRO demonstrates superior caption quality under human evaluation.
-This aligns with the result in Figure 7, which is further supported by the high consistency of 85%.
+For this, 100 questions are randomly sampled from the testmini set of MathVista, and captions are generated using Qwen2.5-VL-3B, trained with and without VPO.
+A total of 4 trained human annotators are recruited, with each annotator comparing all the captions pairs to determine a winner or a tie.
+For each caption pair, we aggregate the results from different annotators by taking the majority of the decisions.
+Specifically, there are 4 annotators and only 3 decisions (win, tie and lose), so there is at least one decision that occurs twice.
+We compute the inter-annotator consistency following Zheng et al.
+(2023) by calculating the ratio of identical decision pairs out of all possible decision pairs and average them across all samples.
+In Table 8, we report the win/tie/lose ratio (i.e., “win” means captions generated by MLLMs with VPO is better) and an additional measure of GPT-human consistency, calculated by the agreement rate between GPT-4o and human judgments.
+As can be seen, Qwen2.5-VL-3B trained with VPO demonstrates superior caption quality under human evaluation.
+This aligns with the result in Figure 11, which is further supported by the high consistency of 87%.
 This supports the rationale of using GPT-4o as a judge for evaluating caption quality.
 Table 8: Human evaluation on pairwise comparison of the caption quality.
-Win Tie Lose Consistency
-60% 29% 11% 85%
-C.3 Training Dynamics of CRO
-We show the average reward scores and caption lengths over training steps in Figure 13. We observe the following:
-  • Rewards increased as training progresses. This confirms the effectiveness of CRO as it allows the MLLM to generate captions that lead to higher reasoning accuracy.
+Win Tie Lose GPT-human consistency Inter-annotator consistency
+62% 32% 6% 87% 85%
+E.4 Training Dynamics of VPO
+We show the average reward and caption lengths over training in Figures 21 and 22. We observe that:
+  • Rewards increased as training progresses. This confirms the effectiveness of VPO as it allows the MLLM to generate captions that lead to higher reasoning accuracy.
   • Caption lengths grow as training progresses.
 An explanation for this phenomenon is that the MLLM learns to generate more comprehensive captions during training, which is reflected by longer lengths.
-This is also confirmed in the Appendix D, where we visualize the captions.
+This is also confirmed in the Appendix G, where we visualize the captions.
 (a) Qwen2.5-VL-3B
@@
 (c) Qwen2.5-VL-32B
-Figure 13: Training dynamics of the CRO process: Reward and Length Over Steps for Qwen2.5-VL-3B, Qwen2.5-VL-7B and Qwen2.5-VL-32B
-Appendix D Case Study
+(d) Qwen2.5-VL-72B
+Figure 21: Rewards over training steps for VPO.
+(a) Qwen2.5-VL-3B
+(b) Qwen2.5-VL-7B
+(c) Qwen2.5-VL-32B
+(d) Qwen2.5-VL-72B
+Figure 22: Caption length over training steps for VPO.
+E.5 Training Dynamics of GRPO
+We visualize the training dynamics of GRPO in Figure 23.
+For 3B and 7B MLLMs, the rewards fluctuate and eventually drop after an initial convergence.
+In contrast, the 72B model exhibits a more stable convergence.
+In both cases, subsequent VPO stage continues to improve the performance under the perception-reasoning pipelines.
+(a) Qwen2.5-VL-3B
+(b) Qwen2.5-VL-7B
+(c) Qwen2.5-VL-72B
+Figure 23: Rewards over training steps for GRPO.
+Appendix F Use of Large Language Models (LLMs)
+In this paper, the role of Large Language Models (LLMs) was confined to a minor, supporting capacity for polishing the writing.
+They were not involved in the core research process, such as ideation or analysis.
+Appendix G Case Study
 Caption qualities.
-We conduct a case study on the generated query-conditioned captions.
+We conduct a case study on the generated query-relevant captions.
 Specifically, for a multi-modal reasoning question and image, we investigate the quality of the generated captions.
-For MLLMs, we consider Qwen2.5-VL series (3B/32B) both with and without CRO.
+For MLLMs, we consider Qwen2.5-VL series (3B/32B) both with and without VPO.
 We visualize the question, image and captions in Tables LABEL:tbl:case_3b_1- LABEL:tbl:case_3b_7 for Qwen2.5-VL-3B and Tables LABEL:tbl:case_32b_1- LABEL:tbl:case_32b_5 for Qwen2.5-VL-32B.
-Comparing the captions generated by MLLMs with and without CRO, we discover the following:
-  • CRO leads to more visual details.
+Comparing the captions generated by MLLMs with and without VPO, we discover the following:
+  • VPO leads to more visual details.
 We highlight these visual details in red in the table.
 Notably, these details are important clues required to correctly solve the question.
-This shows that CRO is effective in improving the quality (especially comprehensiveness) of the query-conditioned captions.
-  • CRO leads to captions with more organized and hierarchical structures.
-For example, in Table LABEL:tbl:case_3b_6, the MLLM with CRO describes the images at three levels, i.e., Tropic level, Terrestrial food chain and aquatic food chain.
+This shows that VPO is effective in improving the quality (especially comprehensiveness) of the query-relevant captions.
+  • VPO leads to captions with more organized and hierarchical structures.
+For example, in Table LABEL:tbl:case_3b_6, the MLLM with VPO describes the images at three levels, i.e., Tropic level, Terrestrial food chain and aquatic food chain.
 This allows the reasoner to quickly locate important information in the captions.
@@
 Reasoning accuracies.
-In Table LABEL:tbl:r_case_3b_1, we provide a complete comparison of captions (generated by Qwen2.5-VL-3B with and without CRO) and the resulting reasoning results (produced by R1-7B).
-Similarly to the case study on caption quality, MLLMs trained with CRO generate captions that capture more details.
+In Table LABEL:tbl:r_case_3b_1, we provide a complete comparison of captions (generated by Qwen2.5-VL-3B with and without VPO) and the resulting reasoning results (produced by R1-7B).
+Similarly to the case study on caption quality, MLLMs trained with VPO generate captions that capture more details.
 This is critical to the correctness of the reasoning process.
-As can be seen, the reasoner that receives captions with CRO arrives at the correct answer after several rounds of thinking and reflection.
+As can be seen, the reasoner that receives captions with VPO arrives at the correct answer after several rounds of thinking and reflection.
 However, the reasoner that accepts the under-optimized caption experiences multiple contradictions and confusion (highlighted in brown), which leads to responses that exceeds the maximum context length and finally fails this problem.
-Table 7: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Appendix H Limitation
+Auto-thinking.
+RAPID is specifically designed for multi-modal reasoning, and it is appealing to explore how to flexibly switch between fast and slow thinking dependent on the complexity of input queries, without human prior and prompt engineering.
+Domain-Specific Design.
+The RAPID architecture, in its current form, is optimized for multi-modal math and science reasoning.
+Its extension to other domains, such as spatial reasoning (Wang et al., 2024a), would require more than re-evaluation; it would demand specific adaptations to the model itself.
+The experiments on general benchmarks reported in Sec.
+4.1 do not test the full reasoning pipeline, as the LLM was not activated.
+Therefore, investigating the adaptations required to generalize RAPID remains a key open direction for future research.
+Adapting the LLM.
+In the current implementation of RAPID, the LLM functions as a static reasoning module with its parameters kept frozen.
+A valuable direction for future work is to explore adapting the LLM to our perception-reasoning framework, potentially through methods like supervised fine-tuning or reinforcement learning.
+While this would introduce additional computational overhead for training, it remains an open empirical question whether the potential performance gains would justify the increased cost.
+Table 9: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 (A)Cannot say (B) 1,500 (C) 125 (D) 150 (E) 1,250
-Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
 The image is a bar graph titled “Fuel Consumption (miles to the gallon)” that compares the fuel efficiency of five different cars, labeled as Taber, Ursa, Velvo, Tink, and Xtam.
@@
 - Cost to Purchase (£): Taber (12,500), Ursa (15,250), Velvo (37,500), Tink (55,250), Xtam (62,500) The image does not contain any information related to the question about fuel consumption for a specific car model.
-Query-conditioned Captions (Qwen2.5-VL-3B):
+Query-relevant Captions (Qwen2.5-VL-3B):
 The image is a bar graph titled “Fuel Consumption (miles to the gallon)” that compares the fuel efficiency of five different cars, Taber, Ursa, Velvo, Tink, and Xtam, under two driving conditions: city driving and motorway driving.
@@
 Cost to Purchase: £62,500 The graph shows that the Taber and Ursa cars have higher fuel efficiency in city driving compared to motorway driving, while the Velvo, Tink, and Xtam cars have higher fuel efficiency in motorway driving compared to city driving.
-Table 8: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 10: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 According to the question shown in the image, please directly answer the question and provide the correct option letter, e.g., A, B, C, D.
-Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
 The image is a geometric diagram with the following elements:
@@
 The question asks for the measure of angle $2$, given that $AB\parallel CD$ and $EG$ bisects angle $\angle BEF$.
-Query-conditioned Captions (Qwen2.5-VL-3B):
+Query-relevant Captions (Qwen2.5-VL-3B):
 The image shows a geometric figure with lines AB and CD parallel to each other. Line EG bisects angle BEF. Angle 1 is marked as 50 degrees.
-Table 9: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 11: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 Input binary number B = 11000, find the corresponding decimal value by the algorithm shown in image.
-Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
 The image is a flowchart titled "Start Binary To Decimal." It outlines a process to convert a binary number to its decimal equivalent.
@@
 The image does not contain any visual elements or additional information beyond the flowchart itself.
-Query-conditioned Captions (Qwen2.5-VL-3B):
+Query-relevant Captions (Qwen2.5-VL-3B):
 Start Binary To Decimal
@@
 Power ++
-Table 10: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 12: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 The pie chart below represents the number of each grade at a United States elementary school. What is the percentage of Grade 2? The answer should be a percentage between 0 and 100.
-Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
 The image is a pie chart titled "Number of Students in Each Grade." The chart is divided into five segments, each representing a different grade level and the number of students in that grade.
@@
 The chart provides a clear visual representation of the number of students in each grade at the elementary school.
-Query-conditioned Captions (Qwen2.5-VL-3B):
+Query-relevant Captions (Qwen2.5-VL-3B):
 The pie chart below represents the number of each grade at a United States elementary school. What is the percentage of Grade 2? The answer should be a percentage between 0 and 100.
-Table 11: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 13: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 Is the left graph isomorphic with the right graph? choice: (A) Yes (B) No
-Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
 The image consists of two graphs.
@@
 The left graph has a more complex structure with multiple connections and a central node (E), while the right graph is a simple cycle with no such central node and a different number of vertices.
-Query-conditioned Captions (Qwen2.5-VL-3B):
+Query-relevant Captions (Qwen2.5-VL-3B):
 The left graph is a graph with 8 vertices and 10 edges. The right graph is a graph with 8 vertices and 8 edges. The left graph is not isomorphic to the right graph.
-Table 12: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 14: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 Question: What happens to fish if pelicans increase?Choices: (A) decrease (B) nothing (C) increase (D) none of the above
-Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
 The image depicts a food chain diagram showing the trophic levels and different types of consumers in both terrestrial and aquatic environments.
@@
 - Tertiary consumer (secondary carnivore): Pelican The image also shows arrows indicating the direction of energy flow from one trophic level to the next, emphasizing the flow of energy through different ecosystems.
-Query-conditioned Captions (Qwen2.5-VL-3B): The image shows a food chain diagram with two types of food chains: terrestrial and aquatic.
+Query-relevant Captions (Qwen2.5-VL-3B): The image shows a food chain diagram with two types of food chains: terrestrial and aquatic.
 The terrestrial food chain includes a caterpillar, lizard, fish, and snake, with the caterpillar being the primary producer and the snake being the tertiary consumer.
 The aquatic food chain includes zooplankton, phytoplankton, fish, and a pelican, with the zooplankton being the primary producer and the pelican being the tertiary consumer.
-Table 13: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 15: Case Study with Qwen2.5-VL-3B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 Choices:(A) crickets (B) deer (C) snakes (D) hawks
-Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
 The image is a food web diagram depicting various organisms and their interactions within an ecosystem.
@@
 The arrows indicate the direction of energy flow from one organism to another, showing the relationships between the different species in the ecosystem.
-Query-conditioned Captions (Qwen2.5-VL-3B): The image is a food web diagram depicting various organisms and their interactions within an ecosystem.
+Query-relevant Captions (Qwen2.5-VL-3B): The image is a food web diagram depicting various organisms and their interactions within an ecosystem.
 It includes a deer, trees, grasses, mice, crickets, frogs, hawks, and owls.
@@
 The diagram shows how energy moves through the ecosystem, with each organism feeding on others, and how the removal of frogs could potentially lead to an increase in the population of crickets, as they are a primary food source for frogs.
-Table 14: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 16: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end. Question: Which option is correct? Choices: (A) A (B) B (C) C
-Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
 The image presents a logic puzzle involving patterns and shapes.
@@
 The internal designs also follow a pattern: Star $\rightarrow$ Square $\rightarrow$ Downward Arrow.
-Query-conditioned Captions (Qwen2.5-VL-32B):
+Query-relevant Captions (Qwen2.5-VL-32B):
 The image consists of two main sections: Left Section: A 3x3 grid with various shapes and patterns.
@@
 The task appears to involve identifying a pattern or sequence based on the shapes and their internal symbols in the grid and selecting the correct option from the choices provided.
-Table 15: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 17: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 Hint: Please answer the question requiring an integer answer and provide the final value, e.g., 1, 2, 3, at the end. Question: How many models in the figure achieve an Acc score greater than 60?
-Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
 The image is a scatter plot titled "Model Performance," which compares the performance of various large language models on a specific task, likely related to the GSM8K dataset.
@@
 Notes that GPT-4 uses some part of the GSM8K data in pretraining and suggests its performance lies between SFT and ICL.
-Query-conditioned Captions (Qwen2.5-VL-32B):
+Query-relevant Captions (Qwen2.5-VL-32B):
 The image is a scatter plot titled "Model Performance." It compares the performance of various large language models on a task, likely related to the GSM8K dataset, based on two metrics: "ACC" (accuracy) on the y-axis and "Pretrain Loss" on the x-axis.
@@
 However, the instruction is to describe the image without answering the question.
-Table 16: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 18: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 Hint: Please answer the question and provide the correct option letter, e.g., A, B, C, D, at the end. Question: Does South Carolina have the highest value in the South ? Choices: (A) Yes (B) No
-Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
 The image is a choropleth map of the United States, depicting the distribution of health insurance coverage for nonelderly individuals (ages 0–64) through employer-sponsored plans in 2013.
@@
 The highest values in the South are represented by Texas and Florida, which are in the darkest blue category.
-Query-conditioned Captions (Qwen2.5-VL-32B):
+Query-relevant Captions (Qwen2.5-VL-32B):
 The image is a choropleth map of the United States, illustrating the distribution of health insurance coverage for nonelderly individuals aged 0–64 through employer-sponsored plans in 2013.
@@
 The map shows South Carolina in a light blue shade, indicating it falls into the second-lowest category of coverage.
-Table 17: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 19: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 When the switch is closed what would happen to the potential difference across the 15 ohm resistor?
-Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
 The image depicts an electrical circuit involving a battery, an ammeter, three resistors, and a switch.
@@
 Therefore, the potential difference across the 15 $\Omega$ resistor is equal to that across the 30 $\Omega$ resistor.
-Query-conditioned Captions (Qwen2.5-VL-32B):
+Query-relevant Captions (Qwen2.5-VL-32B):
 The image depicts a simple electrical circuit consisting of the following components: Battery: A power source labeled as "Battery" is shown on the left side of the circuit.
@@
 This description provides a detailed overview of the circuit’s components and their arrangement.
-Table 18: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Table 20: Case Study with Qwen2.5-VL-32B. Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Image:
@@
 Choices: (A) A (B) B (C) C (D) D (E) E
-Query-conditioned Captions (Qwen2.5-VL-32B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-32B w/ VPO):
 The image presents a multiple-choice question involving illustrations of cats.
@@
 Section D: 2 striped, 1 spotty, 2 white kittens. One kitten has ears of a different color. Section E: 1 striped, 3 spotty, 1 white kittens. One kitten has ears of a different color.
-Query-conditioned Captions (Qwen2.5-VL-32B):
-[…]
+Query-relevant Captions (Qwen2.5-VL-32B):
+The image shows five groups of cartoon-style cats, each labeled with a letter (A, B, C, D, E).
+Each group contains multiple cats with various patterns and colors, including striped, spotty, and solid-colored cats.
+The cats are depicted in different poses, such as sitting, lying down, or standing.
+The question asks to identify the group that matches the description of Misty’s kittens, which includes two striped, one spotty, and two white kittens, with one having differently colored ears.
+Detailed Description of Each Group: Group A: Contains cats with striped and spotty patterns.
+Some cats have differently colored ears.
+Group B: Features cats with striped and solid-colored patterns.
+No cats have differently colored ears.
+Group C: Includes cats with striped, spotty, and solid-colored patterns.
+No cats have differently colored ears.
+Group D: Shows cats with striped and solid-colored patterns.
+No cats have differently colored ears.
+Group E: Contains cats with striped, spotty, and solid-colored patterns.
+No cats have differently colored ears.
+The task is to identify the group that matches the specific description of Misty’s kittens.
+Misty’s Kittens: Two striped kittens One spotty kitten Two white (solid-colored) kittens One kitten has differently colored ears Conclusion: Only Group A meets all the criteria, including the presence of differently colored ears.
+$\boxed{\text{A}}$
 Table 21: Reasoning Case Study with Qwen2.5-VL-3B.
-Additional visual details that are relevant to the question but are only generated by MLLMs with CRO are highlighted in red.
+Additional visual details that are relevant to the question but are only generated by MLLMs with VPO are highlighted in red.
 Reasoning process that show confusing and contradiction are highlighted in brown.
@@
 If point $D$ is the midpoint of $BC$, what is the length of $AC$?
-Query-conditioned Captions (Qwen2.5-VL-3B w/ CRO):
+Query-relevant Captions (Qwen2.5-VL-3B w/ VPO):
 The image is a geometric diagram of a triangle $\triangle ABC$ inscribed in a circle with center $O$. The diagram shows the following elements:
@@
 The focus is purely on the geometric relationships and the diagram itself.
-Query-conditioned Captions (Qwen2.5-VL-3B):
+Query-relevant Captions (Qwen2.5-VL-3B):
 The image shows a circle with center $O$ and points $A$, $B$, and $C$ on its circumference.
@@
 Point $D$ is the midpoint of $BC$.
-Reasoning Result of R1-7B (captions from Qwen2.5-VL-3B w/ CRO):
+Reasoning Result of R1-7B (captions from Qwen2.5-VL-3B w/ VPO):
 Alright, let me try to figure out this geometry problem step by step.
 So, we have triangle $ABC$ inscribed in a circle with center $O$.
-$AB$ and $BC$ are both 16 units long, making it an isosceles triangle with $AB=BC$.
+ABAB and $BC$ are both 16 units long, making it an isosceles triangle with $AB=BC$.
 The diagram shows that when arc $AB$ is folded along chord $AB$, it intersects $BC$ at point $D$, which is the midpoint of $BC$.
@@
 So, AB = AC = 16. … exceeding maximum context length (8192)
-Appendix E Broader Impacts
-RACRO improves multi-modal reasoning by aligning perception with reasoning through reinforcement learning.
-While beneficial for tasks requiring precise vision-language understanding, this alignment could be misused for generating misleading content or surveillance.
-Caution is advised when applying RACRO in sensitive domains.
-Mitigation strategies include human oversight, responsible deployment, and transparency in model behavior.
-Appendix F Limitation
-We evaluate the effectiveness of the proposed RACRO framework exclusively on multi-modal reasoning tasks.
-However, its potential applicability could extend to other important domains, such as medical visual question answering (VQA) and geospatial reasoning.
-Exploring these areas represents promising directions for future work.
-Appendix G More Discussions
-Future works.
-This work aims at decoupling multi-modal reasoning with visual perception and LLM reasoning, and the successful visual perception is essential for downstream reasoning.
-How to further enhance the visual perception capability is a prevailing research direction (e.g., utilizing AIGC-generated data [10, 16, 17, 18, 23, 28, 29, 33, 50] and enhanced visual representations [7, 8, 35, 58])